### v1

# On the Kantorovich contraction of Markov semigroups

Abstract
This paper develops a novel operator theoretic framework to study the contraction properties of Markov semigroups with respect to a general class of Kantorovich semi-distances, which notably includes Wasserstein distances. The rather simple contraction cost framework developed in this article, which combines standard Lyapunov techniques with local contraction conditions, helps to unifying and simplifying many arguments in the stability of Markov semigroups, as well as to improve upon some existing results.
Our results can be applied to both discrete time and continuous time Markov semigroups, and we illustrate their wide applicability in the context of (i) Markov transitions on models with boundary states, including bounded domains with entrance boundaries, (ii) operator products of a Markov kernel and its adjoint, including two-block-type Gibbs samplers, (iii) iterated random functions and (iv) diffusion models, including overdampted Langevin diffusion with convex at infinity potentials.
Keywords: Markov operators and semigroups, contraction coefficients, Kantorovich semi-distances, Wasserstein distances, weighted total variation norms.
Mathematics Subject Classification: Primary 37H30, 60J05, 47H09; secondary 37A30, 60J25.
1 Introduction
1.1 Set-up
In this work we let ${\cal B}(S)$ denote the set of measurable functions on a complete separable metric space $(S,\psi_{S})$, we let $\|f\|:=\sup_{x\in S}|f(x)|$ denote the uniform norm of a function $f\in{\cal B}(S)$ and ${\cal P}(S)$ be the set of probability measures on $S$. For any measures $\mu_{1},\mu_{2}\in{\cal P}(S)$ we let $\Pi(\mu_{1},\mu_{2})$ denote the convex subset of probability measures $\pi\in{\cal P}(S^{2})$ with respective marginal $\mu_{1}$ and $\mu_{2}$ w.r.t. the first and the second coordinate.
Next, we let $V\in{\cal B}(S)$ be a lower semi-continuous (abbreviated l.s.c.) function, lower bounded away from zero, and we denote by ${\cal B}_{V}(S)$ the subspace of functions $f\in{\cal B}(S)$ such that $\|f/V\|<\infty$. In addition, we set
${\cal P}_{V}(S):=\left\{\mu\in{\cal P}(S)~:~\mu(V)<\infty\right\}\quad\mbox{\rm with}\quad\mu(V):=\int~V(x)~\mu(dx)$
and recall that the weighted total variation norm (a.k.a. $V$-norm) on ${\cal P}_{V}(S)$ is defined, for any $\mu_{1},\mu_{2}\in{\cal P}_{V}(S)$, by the formula
$\|\mu_{1}-\mu_{2}\|_{V}:=\sup{\left\{|(\mu_{1}-\mu_{2})(f)|~:~f\in{\cal B}_{V}(S)~s.t.~\|f/V\|\leq 1\right\}}=|\mu_{1}-\mu_{2}|(V).$
In the above display $|\nu|:=\nu_{+}+\nu_{-}$ stands for the total variation measure associated with a Hahn-Jordan decomposition $\nu=\nu_{+}-\nu_{-}$ of the signed measure $\nu=\mu_{1}-\mu_{2}$.
The right-action $f\in{\cal B}_{V}(S)\mapsto P(f)\in{\cal B}_{V}(S)$ and the left-action $\mu\in{\cal P}_{V}(S)\mapsto\mu P\in{\cal P}_{V}(S)$ of a Markov integral operator $P(x,dy)$ are defined, for any $x\in S$, by the function
$P(f)(x)=\int~P(x,dy)~f(y)\quad\mbox{\rm and the measure}\quad(\mu P)(dy):=\int\mu(dx)P(x,dy).$ (1.1)
In what follows, we say that $P$ is a $V$-operator as soon as we have
${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|P\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\tiny op,V}:=\|P(V)/V\|=\sup{\left\{\|P(f)/V\|~:~f\in{\cal B}_{V}(S)~\mbox{s.t.}~\|f/V\|\leq 1\right\}}<\infty$ (1.2)
and we denote by $(P_{n})_{n\geq 0}$ the discrete generation semigroup associated with $P$, defined sequentially for any integer $n\geq 1$ by
$P_{n}:=P_{n-1}P=PP_{n-1}\quad\text{with $P_{0}=I$ the identity operator}.$ (1.3)
In the above display, $P_{n-1}P$ is a shorthand notation for the composition $P_{n-1}\circ P$ of the right or left-action operators defined respectively in (1.1).
We say that a function $\phi$ defined on $S^{2}$ is a semi-distance if the following three conditions hold: (i) $\phi(x,y)\geq 0$ for all $x,y\in S$, (ii) $\phi(x,y)=0$ if and only if $x=y$ and (iii) $\phi$ is l.s.c. (and therefore measurable). It is easily checked that semi-distances are stable w.r.t. sums and products, and that $h\circ\phi$ is a semi-distance if $h$ is strictly increasing with $h(0)=0$ and $\phi$ is a semi-distance.
Examples of semi-distances discussed in the article include the discrete metric $\varphi_{0}(x,y):=1_{x\neq y}$ and, since $V$ is assumed to be l.s.c., the weighted discrete metric defined by
$\varphi_{V}(x,y):=\varphi_{0}(x,y)~\varpi_{V}(x,y)\quad\mbox{\rm with}\quad\varpi_{V}(x,y)=V(x)+V(y).$ (1.4)
Note also that for any semi-distances $(\phi,\kappa)$ and any $\iota\in[0,1]$ the functions
$\phi_{V}(x,y):=\phi(x,y)+\varphi_{V}(x,y)\quad\mbox{\rm and}\quad\kappa_{\iota,V}(x,y):=\kappa(x,y)^{\iota}~\varphi_{V}(x,y)^{1-\iota}$ (1.5)
are also semi-distances.
We end this section with some notation and terminology used in the article. For indicator functions $f=1_{A}$ of a measurable subset $A\subset S$ and a given measure $\nu$ sometimes we write $\nu(A)$ instead of $\nu(1_{A})$. A measure $\nu_{1}$ is said to be absolutely continuous with respect to another measure $\nu_{2}$, and we write $\nu_{1}\ll\nu_{2}$, if $\nu_{2}(A)=0$ implies $\nu_{1}(A)=0$ for any measurable subset $A\subset S$. Whenever $\nu_{1}\ll\nu_{2}$, we denote by ${d\nu_{1}}/{d\nu_{2}}$ the Radon-Nikodym derivative of $\nu_{1}$ w.r.t. $\nu_{2}$. We also consider the partial order, denoted $\nu_{1}\leq\nu_{2}$, when $\nu_{1}(A)\leq\nu_{2}(A)$ for any measurable subset $A\subset S$. The positive and negative part of $a\in\mathbb{R}$ are defined respectively by $a_{+}=\max(a,0)$ and $a_{-}=\max(-a,0)$. Given $a,b\in\mathbb{R}$ we also set $a\vee b=\max(a,b)$ and $a\wedge b=\min(a,b)$.
1.2 Kantorovich semi-distances
For some given semi-distance $\phi$, the Kantorovich semi-distance between two measures $\mu_{1},\mu_{2}\in\mathcal{P}(S)$ is defined by
$D_{\phi}(\mu_{1},\mu_{2}):=\inf_{\pi\in\Pi(\mu_{1},\mu_{2})}{\pi(\phi)}\quad\mbox{\rm with}\quad\pi(\phi):=\int~\phi(x,y)~\pi(d(x,y)).$ (1.6)
The assumption that $S$ is a complete and separable metric space and the fact that $\phi$ is l.s.c. and non-negative ensure the existence of an optimal transport plan (see [72], Theorem 4.1). Moreover, the assumption that $\phi(x,y)=0\Leftrightarrow x=y$ ensure that $D_{\phi}(\mu_{1},\mu_{2})=0$ if and only if $\mu_{1}=\mu_{2}$.
Noting that $D_{\phi}(\delta_{x},\delta_{y})=\phi(x,y)$, the quantity $\phi(x,y)$ can be seen as measuring the cost of transporting $x$ to $y\in S$, and it is typically small when the states $x$ and $y$ are close. We also note that in the optimal transport literature, the measures in $\Pi(\mu_{1},\mu_{2})$ are sometimes called Kantorovich plans, transport plans or couplings.
The Kantorovich semi-distance is related to several well-known measures of distance between probability distributions. Specifically, the quantity $D_{\psi_{S}}(\mu_{1},\mu_{2})$ is the Kantorovitch distance between $\mu_{1}$ and $\mu_{2}$ (a.k.a. the Wasserstein 1-distance) while for $p\geq 1$ the Wasserstein $p$-distance $\mathbb{W}_{\psi_{S},p}(\mu_{1},\mu_{2})$ is defined by $\mathbb{W}_{\psi_{S},p}(\mu_{1},\mu_{2})=D_{\psi_{S}^{p}}(\mu_{1},\mu_{2})^{p}$. We also have the equality $D_{\varphi_{0}}(\mu_{1},\mu_{2})=\|\mu_{1}-\mu_{2}\|_{tv}$, that is $D_{\varphi_{0}}(\mu_{1},\mu_{2})$ is the total variation distance between the probability measures $\mu_{1}$ and $\mu_{2}$, and the weighted total variation norm $\|\mu_{1}-\mu_{2}\|_{V}$ is related to the Kantorovich semi-distance (1.6) by the formula
$D_{\varphi_{V}}(\mu_{1},\mu_{2})=\|\mu_{1}-\mu_{2}\|_{V}.$ (1.7)
We also have the dual formulation
$\|\mu_{1}-\mu_{2}\|_{V}:=\sup{\left\{|(\mu_{1}-\mu_{2})(f)|~:~f\in{\cal B}(S)~;~\mbox{\rm osc}_{V}(f)\leq 1\right\}},$ (1.8)
with the $\varphi_{V}$-oscillation $\mbox{\rm osc}_{V}(f)$ of the function $f\in{\cal B}(S)$ defined by the Lipschitz constant
$\mbox{\rm osc}_{V}(f):=\sup{\left\{|f(x_{1})-f(x_{2})|/\varphi_{V}(x_{1},x_{2})~:~(x_{1},x_{2})\in S^{2}~;~x_{1}\not=x_{2}\right\}}.$
This $V$-norm formulation of the Kantorovich semi-distance associated with the weighted discrete metric $\varphi_{V}$ is a consequence of the Kantorovich-Rubinstein theorem [33, 50, 51]; see Theorem 19.1.7 in [28] and also Proposition 8.2.16 in [22] for a proof of (1.7) and (1.8).
If the assumption that $\phi$ is a semi-distance ensures the existence of an optimal transport plan, it does not guarantee that the quantity $D_{\phi}(\mu_{1},\mu_{2})$ is finite. To solve this issue we assume that $\phi$ is a $V$-semi-distance, in the sense that $\|\phi/\varpi_{V}\|<\infty$. Under this assumption on $\phi$, for any $\mu_{1},\mu_{2}\in{\cal P}_{V}(S)$ we indeed have
$D_{\phi}(\mu_{1},\mu_{2})\leq(\mu_{1}\otimes\mu_{2})(\phi)\leq\mu_{1}(V)+\mu_{2}(V)<\infty$
with the tensor product measure $(\mu_{1}\otimes\mu_{2})(d(x,y)):=\mu_{1}(dx)\mu_{2}(dy)$.
1.3 Contraction of Markov transport maps
By successively applying a Markov operator $P$ to the probability measures $\mu_{1}$ and $\mu_{2}$ we expect the measures $\mu_{1}P_{n}$ and $\mu_{2}P_{n}$ to become closer as $n$ increases, so that the Kantorovitch semi-distance between $\mu_{1}P_{n}$ and $\mu_{2}P_{n}$ decreases with $n$.
To quantify this property, for two $V$-semi distances $\phi$ and $\psi$ and for a Markov $V$-operator $P$ we let $\beta_{\psi,\phi}(P)\in[0,\infty]$ be defined by
$\beta_{\psi,\phi}(P)=\sup{\frac{D_{\phi}(\mu_{1}P,\mu_{2}P)}{D_{\psi}(\mu_{1},\mu_{2})}}$ (1.9)
where the supremum is taken over all pairs of measures $(\mu_{1},\mu_{2})\in{\cal P}_{V}(S)^{2}$ for which we have $D_{\psi}(\mu_{1},\mu_{2})>0$. Notice that, since $P$ is a Markov $V$-operator and $\mu_{1},\mu_{2}\in\mathcal{P}_{V}(S)$, both $\mu_{1}P$ and $\mu_{2}P$ belong to $\mathcal{P}_{V}(S)$ and thus, following the discussion of the previous subsection, $D_{\phi}(\mu_{1}P,\mu_{2}P)<\infty$; we however stress that the coefficient defined in (1.9) may not be finite.
This coefficient, which we will refer to as the $(\psi,\phi)$-Dobrushin contraction coefficient of $P$ in what follows, does not depend on the function $V$, and the supremum in (1.9) is obtained on Dirac measures. Specifically, for any $V$-semi-distances $\phi$ and $\psi$ we let $\ell_{\psi,\phi}(P):(x,y)\in S^{2}\mapsto[0,\infty]$ be the function defined by
$\ell_{\psi,\phi}(P)(x,y)=\frac{D_{\phi}(\delta_{x}P,\delta_{y}P)}{D_{\psi}(\delta_{x},\delta_{y})},\quad x,y\in S$
and prove the following result:
Lemma 1.1.
Let $\phi$ and $\psi$ be two $V$-semi-distances. Then,
$\beta_{\psi,\phi}(P)=\sup_{(x,y)\in S^{2}~:~x\neq y}\ell_{\psi,\phi}(P)(x,y).$ (1.10)
Proof.
See Section 4.1. ∎
Given a $V$-semi-distance $\phi$, a first natural choice for the reference $V$-semi-distance $\psi$ used to define the coefficient in (1.9) is $\psi=\phi$. In this case the coefficient $\beta_{\phi}(P):=\beta_{\phi,\phi}(P)$ is the so-called $\phi$-Dobrushin contraction coefficient (a.k.a. Kantorovich norm [27]), and below we will write $\ell_{\phi}(P)$ instead of $\ell_{\phi,\phi}(P)$. In particular, it is worth noting that, by (1.7), the coefficient $\beta_{\varphi_{V}}(P)$ is equal to the $V$-norm contraction coefficient of $P$. A second natural choice for the $V$-semi-distance $\psi$ in (1.9) is the $V$-semi-distance $\psi=\varphi_{V}$, in which case we have, by (1.7),
$\beta_{\varphi_{V},\phi}(P)=\sup\frac{D_{\phi}(\mu_{1}P,\mu_{2}P)}{\|\mu_{1}-\mu_{2}\|_{V}},$
where the supremum is taken over all measures $(\mu_{1},\mu_{2})\in{\cal P}_{V}(S)^{2}$ such that $\|\mu_{1}-\mu_{2}\|_{V}>0$.
Consider a Markov integral operator $K$ and some other l.s.c. function $W$, lower bounded away from zero and such that $\|K(W)/V\|<\infty$. In this case, the right-action $f\in{\cal B}_{W}(S)\mapsto K(f)\in{\cal B}_{V}(S)$ and the left-action $\mu\in{\cal P}_{V}(S)\mapsto\mu K\in{\cal P}_{W}(S)$ are well defined and, by (1.7), we have
$\beta_{\varphi_{V},\varphi_{W}}(K)=\raisebox{2.0pt}{$\chi$}_{V,W}(K):=\sup\frac{\|\mu_{1}K-\mu_{2}K\|_{W}}{\|\mu_{1}-\mu_{2}\|_{V}}$ (1.11)
where the supremum is taken over all measures $(\mu_{1},\mu_{2})\in{\cal P}_{V}(S)^{2}$ such that $\|\mu_{1}-\mu_{2}\|_{V}>0$. The condition $\|K(W)/V\|<\infty$ ensures that $K$ is a $(V,W)$-operator in the sense that
${\left|\kern-1.07639pt\left|\kern-1.07639pt\left|K\right|\kern-1.07639pt\right|\kern-1.07639pt\right|}_{\tiny V,W}:=\|K(W)/V\|=\sup{\left\{\|K(f)/V\|~:~f\in{\cal B}_{W}(S)~;~\|f/W\|\leq 1\right\}}<\infty$ (1.12)
and, in this context, by the dual formulation (1.8) we have the equivalent formulation
$\raisebox{2.0pt}{$\chi$}_{V,W}(K)=\sup{\{\mbox{\rm osc}_{V}(K(f))~:~f\in{\cal B}(S)~;~\mbox{\rm osc}_{W}(f)\leq 1\}}.$ (1.13)
For a more detailed discussion on these equivalent formulations we refer to Section 8.2 in [22] and, for completeness, a sketched proof of (1.13) is provided in Section 4.2.
Lastly, we note that the product $P:=KL$ of a $(V,W)$-operator $K$ and a $(W,V)$-operator $L$ is a $V$-operator, and we have
[…]
1.4 Contribution of the paper
The purpose of this article is to provide conditions on the pair $(P,\phi)$ of Markov $V$-operator and $V$-semi distance under which the contraction coefficients $\beta_{\phi}(P_{n})$ and $\beta_{\varphi_{V},\phi}(P_{n})$ converge to zero exponentially fast as $n\rightarrow\infty$. Our operator theoretic approach extends the one developed in [4, 22] for the weighted total variation norm to a broader class of Kantorovitch semi-distances, which notably includes Wasserstein distances. Our analysis relies on two standard assumptions on $P$; namely a simple geometric drift condition and a local contraction property. Freed from couplings arguments and carefully designed distance-like functions, the proofs of our main theorems have the advantage to be relatively straightforward. At the same time, these theorems allow to recover several known results, and to improve upon others (as discussed in Section 2.7). The present article notably complements the studies [4, 21] dedicated to the $V$-norm stability of positive semigroups. Specifically, Markov semigroups being particular instances of positive semigroups, the Kantorovich contraction analysis developed in the present article applies directly to some classes of semigroups discussed in [4, 21].
We illustrate our results in the context of Markov transitions on bounded domains with entrance boundaries (i.e. boundary states that cannot be reached from the inside), as well as Gibbs samplers and more standard random iterated functions. To the best of our knowledge, the Wasserstein contraction properties we obtain for Markov transitions on domains with boundaries and the ones we obtain for two-block Gibbs samplers and for a general class of iterated random functions are the first results of this type for these classes of models. We also show that our results can be used to easily derive exponential contraction estimates for continuous time Markov semigroups, and illustrate this application of our operator theoretic approach in the context of diffusion models.
The remainder of this article is structured as follows: Our main stability results as well as a review of the existing literature on the contraction properties of Markov operators are presented in Section 2. Section 3 is dedicated to a series of illustrations, and most of the proofs are collected in Section 4.
2 Contraction theorems
2.1 Assumptions
In this section we let $P$ be a Markov $V$-operator for some l.s.c. function $V$, lower bounded away from zero, and all the results presented below will impose the following regularity condition on $P$:
  1. (H1)
    There exist constants $\epsilon\in]0,1[$ and $c\geq 0$ such that $P(V)\leq\epsilon V+c$.
[…]
In many applications $P=KL$ for some Markov integral operators $K$ and $L$ (see Section 3.3). In this context, if there exists a l.s.c. function $W$, lower bounded away from zero and such that
$K(W)\leq\epsilon~V+c\quad\mbox{\rm and}\quad L(V)\leq\epsilon~W+c$ (2.1)
for some constants $\epsilon_{0}\in]0,1[$ and $c_{0}\geq 0$, then $P$ is a Markov $V$-operator and (H1) holds with $\epsilon=\epsilon_{0}^{2}$ and $c=(1+\epsilon_{0})c_{0}$.
In addition to (H1), the results presented in this section will assume that the following additional condition on $P$ is satisfied:
  1. (H2)
    There exist a $V$-semi-distance $\kappa$ such that, for some parameter $r_{0}>0$ and map $\alpha:[r_{0},\infty[\mapsto\alpha(r)\in]0,1[$, we have
    $\varpi_{V}(x,y)\leq r\implies D_{\kappa}\left(\delta_{x}P,\delta_{y}P\right)\leq(1-\alpha(r))~\kappa(x,y),\quad\forall r\geq r_{0}.$
Like (H1), condition (H2) is very standard in the literature (see Section 2.7). In particular, assuming that (H2) holds for the discrete metric (that is with $\kappa=\varphi_{0}$) is equivalent to assuming the following standard local minorization condition:
$\text{$\forall r\geq r_{0}$, $\exists\alpha(r)\in]0,1[$ and $\exists\nu_{r}\in\mathcal{P}(S)$ such that }V(x)\leq r\implies\delta_{x}P\geq\alpha(r)\,\nu_{r}.$ (2.2)
In this scenario, when the function $V$ is bounded (H2) is a minorization condition on the whole state space, which takes the form
$\forall x\in S\qquad\delta_{x}P\geq\jmath\,\nu\quad\mbox{\rm for some $\jmath>0$ and some probability measure $\nu\in\mathcal{P}(S)$. }$
In the reverse angle, when $V\in{\cal B}_{\infty}(S)$ condition (H2) is a local contraction condition and (2.2) is a local minorization condition on the compact sub-level sets of the function $V$. In particular, when $V\in{\cal B}_{\infty}(S)$ the following lemma shows that condition (H2) holds for the semi-distance $\kappa=\varphi_{0}$ under mild assumptions on $P$.
Lemma 2.1.
Assume that $V\in{\cal B}_{\infty}(S)$ and that $P(x,dy)=p(x,y)\eta(dy)$ for some reference measure $\eta$ on $S$ and continuous function $p:S^{2}\rightarrow]0,\infty[$. Then, condition (H2) holds with the semi-distance $\kappa=\varphi_{0}$.
Proof.
For all $r\geq 1$ let $A_{r}=\{x\in S:V(x)\leq r\}$. Then, since $V\in{\cal B}_{\infty}(S)$ the set $A_{r}$ is compact for all $r\geq 1$ and we have $S=\cup_{n\in\mathbb{N}}A_{n}$, showing that $S$ is locally compact. In this context, there exists an $n^{\prime}\in\mathbb{N}$ such that $\eta(A_{n^{\prime}})>0$, since otherwise $\eta(S)=0$. In addition, since the density $p$ is assumed to be strictly positive and continuous, we have $a_{r}:=\inf_{(x,y)\in A_{r}\times A_{n^{\prime}}}p(x,y)>0$ for all $r\geq n^{\prime}$. Therefore, for all $r\geq n^{\prime}$,
$P(x,dy)\geq P(x,dy)1_{A_{n^{\prime}}}(y)\geq a_{r}\,\eta(dy)1_{A_{n^{\prime}}}(y)=b_{r}q(dy),\quad\forall x\in A_{r}$
with $b_{r}=a_{r}\eta(A_{n^{\prime}})>0$ and $q(dy)=\eta(dy)1_{A_{n^{\prime}}}(y)/\eta(A_{n^{\prime}})$. This shows that (2.2) holds with $r_{0}=n^{\prime}$, $\alpha(r)=(1/2)\wedge b_{r}$ and $\nu_{r}=q$. The proof of the lemma is complete. ∎
2.2 Additional notation
For any $\rho\geq 0$ we introduce the rescaled Lyapunov function $V_{\rho}$ and the l.s.c. function $\varpi_{\rho}$ defined by
$V_{\rho}(x):=\frac{1}{2}+\rho~V(x)\quad\mbox{\rm and}\quad\varpi_{\rho}(x,y):=1+\rho~\varpi_{V}(x,y).$ (2.3)
For any $V$-semi-distances $(\phi,\kappa)$ and any $\iota\in[0,1]$ we also consider the $V$-semi-distances
$\varphi_{\rho}(x,y)$ $:=$ $\varphi_{0}(x,y)~\varpi_{\rho}(x,y)$
$\phi_{\rho}(x,y)$ $:=$ $\phi(x,y)+\varphi_{\rho}(x,y)\quad\mbox{\rm and}\quad\kappa_{\iota,\rho}(x,y):=\kappa(x,y)^{\iota}~\varpi_{\rho}(x,y)^{1-\iota}.$ (2.4)
Note that the functions $(\varpi_{\rho},\varphi_{\rho},\phi_{\rho},\kappa_{\iota,\rho})$ coincide with the functions $(\varpi_{V_{\rho}},\varphi_{V_{\rho}},\phi_{V_{\rho}},\kappa_{\iota,V_{\rho}})$ defined as in (1.4) and (1.5) by replacing $V$ by the rescaled Lyapunov function $V_{\rho}$.
We remark that the l.s.c. functions $\varpi_{\rho}$ and $\varpi_{V}$ are related through the following formula:
$\begin{array}[]{l}\rho\,\varpi_{V}\leq\varpi_{\rho}\leq(1+\rho)\,\varpi_{V},\quad\forall\rho\geq 0.\end{array}$ (2.5)
Lastly, when (H1) and (H2) are met, we set $r_{\epsilon}=1/(1-\epsilon)$ and for any $r>r_{0}\vee r_{\epsilon}$ we define the parameters
$\delta_{\epsilon}(r):=\frac{1-\epsilon}{2+\epsilon}~\left(1-\frac{r_{\epsilon}}{r}\right)\quad\mbox{and}\quad\rho_{\epsilon}(r):=\frac{1}{1+\epsilon}~\frac{\alpha(r)}{2r}~\in~]0,1/2].$ (2.6)
2.3 Some comparison principles
Let $(\phi,\varphi,\psi)$ be three $V$-semi-distances. Then, if $P=KL$ for some Markov $V$-operators $K$ and $L$, we have
$\beta_{\phi,\psi}(P)\leq\beta_{\phi,\varphi}(K)~\beta_{\varphi,\psi}(L).$ (2.7)
Next, we note that
$\|\psi/\varphi\|<\infty\implies\beta_{\phi,\psi}(P)\leq\|\psi/\varphi\|~\beta_{\phi,\varphi}(P)\,\,$
while
$\|\varphi/\phi\|<\infty\implies\beta_{\phi,\psi}(P)\leq~\|\varphi/\phi\|~\beta_{\varphi,\psi}(P).$ (2.8)
Similarly, when $\|\psi/\phi\|<\infty$ we have both
$\beta_{\phi,\psi}(P)\leq\|\psi/\phi\|~\beta_{\phi}(P)\quad\mbox{\rm and}\quad\beta_{\phi,\psi}(P)\leq\|\psi/\phi\|~\beta_{\psi}(P).$
In particular, by applying the latter two results with $\phi=\varphi_{V}$, we have the implication
$\|\psi/\varphi_{V}\|<\infty\Longrightarrow\beta_{\varphi_{V},\psi}(P)\leq\|\psi/\varphi_{V}\|~\beta_{\varphi_{V}}(P)\quad\mbox{\rm and}\quad\beta_{\varphi_{V},\psi}(P)\leq\|\psi/\varphi_{V}\|~\beta_{\psi}(P).$
In the same vein, one can check that
$\|\varphi/\psi\|<\infty\Longrightarrow\beta_{\varphi}(P)\leq\|\varphi/\psi\|~\beta_{\varphi,\psi}(P)\quad\mbox{\rm and}\quad\beta_{\psi}(P)\leq\|\varphi/\psi\|~\beta_{\varphi,\psi}(P).$
The following technical lemma provides some general scaling properties and transfer principles of the $\phi$-Dobrushin coefficient.
Lemma 2.2.
Let $\varphi$ be a $V$-semi-distance function. Then,
  • For any constants $a>0$ and $\iota\in]0,1]$ we have
    $\beta_{a\varphi}(P)=\beta_{\varphi}(P)\quad\mbox{and}\quad\beta_{\varphi^{\iota}}(P)\leq\beta_{\varphi}(P)^{\iota}.$ (2.9)
  • For any $V$-semi-distance functions $\psi$, constants $0<a\leq b$ and $\mu_{1},\mu_{2}\in{\cal P}_{V}(S)$ we have
    $a\varphi\leq\psi\leq b\varphi$ $\Longrightarrow$ $\beta_{\varphi}(P)\leq(b/a)~\beta_{\psi}(P).$ (2.10)
  • For any $V$-semi-distance functions $\varphi\leq\psi$ such that $\beta_{\psi}(P)<1$ and any constants $\iota\geq 0$ and $\delta\in]0,1[$ we have
    $\beta_{\psi}(P)<\frac{\delta}{1+\iota}\Longrightarrow\beta_{\iota\varphi+\psi}(P)\leq\delta.$ (2.11)
    In particular, for $\iota\leq 1-\beta_{\psi}(P)$ we have
    $\beta_{\iota\varphi+\psi}(P)\leq 1-\left(1-\beta_{\psi}(P)\right)^{2}.$ (2.12)
Proof.
See Section 4.3. ∎
2.4 A strict contraction theorem
Theorem 2.3 is the first contraction result of the article. It is expressed in terms of the semi-distances $(\varphi_{\rho},\phi_{\rho},\kappa_{\iota,\rho})$ and the parameters $(\delta_{\epsilon}(r),\rho_{\epsilon}(r))$ defined in (2.4) and (2.6).
For notational convenience it assumes hat (H1) is met with $c=1/2$, but this condition is not restrictive. Indeed, if (H1) holds with $c>1/2$ then the function $\overline{V}:=1/2+\big(\epsilon/(2c)\big)V$ satisfies (H1) with $c=1/2$. Moreover, like $V$, the function $\overline{V}$ is l.s.c. and bounded below away from zero, and $\phi$ is a $V$-semi-distance if and only if $\phi$ is a $\overline{V}$-semi-distance.
Theorem 2.3.
Assume that (H1) is met with $c=1/2$. Then,
  • if (H2) is met for some $V$-semi-distance $\kappa$, for any $r>r_{\epsilon}\vee r_{0}$ and $\iota\in[1/2,1[$ we have the contraction estimate
    $\beta_{\kappa_{\iota,\rho_{\epsilon}(r)}}(P)\leq\Big(1-\alpha(r)\min\big\{\delta_{\epsilon}(r)/2,\alpha(r)\big\}\Big)^{1-\iota}.$ (2.13)
  • if (H2) is met for the discrete distance $\kappa=\varphi_{0}$, for any $r>r_{\epsilon}\vee r_{0}$ we have the contraction estimate
    $\beta_{\varphi_{\rho_{\epsilon}(r)}}(P)\leq 1-\delta_{\epsilon}(r)~\frac{\alpha(r)}{2}.$ (2.14)
    Moreover, for any $V$-semi-distance $\phi$ such that $\|\phi/\varphi_{V}\|\leq\rho_{\epsilon}(r)\big(1-\beta_{\varphi_{\rho_{\epsilon}(r)}}(P)\big)$, we have the contraction estimate
    $\beta_{\phi_{\rho_{\epsilon}(r)}}(P)\leq 1-\left(1-\beta_{\varphi_{\rho_{\epsilon}(r)}}(P)\right)^{2}.$ (2.15)
Proof.
See Section 4.4
The contraction estimate for the weighted discrete metric $\varphi_{\rho}$ stated in (2.14) is a slight variation of a contraction estimate presented in [4] (itself based on the weighted norm contraction analysis presented in [21, 22], see also Lemma 2.5 in [4]). To the best of our knowledge the other contraction estimates presented in Theorem 2.3 are new.
A first merit of the strict contraction estimates given in Theorem 2.3 is to allow to easily show that $P$ has a single invariant distribution $\mu=\mu P\in\mathcal{P}_{V}(S)$. Before stating the next result we recall that $(S,\psi_{S})$ is assumed to be a complete separable metric space.
Corollary 2.4.
Assume that (H1) holds and that (H2) is met either for the $V$-semi-distance $\kappa=\varphi_{0}$ or for some $V$-semi-distance $\kappa$ such that $\|\psi_{S}^{p}/\kappa\|<\infty$ for some $p\geq 1$. Then, $P$ admits a unique invariant distribution $\pi\in\mathcal{P}_{V}(S)$.
Proof.
The fact that the contraction estimate (2.14) implies that $P$ has a single invariant distribution is known (see Theorem 8.2.21 in [22]), and thus below we only prove the second part of the corollary. To simplify notation in what follows we let $\psi=\psi_{S}$. In addition, following the remark made at the of this subsection, to prove the corollary we can without loss of generality assume that (H1) holds with $c=1/2$.
Let $\mathcal{P}_{\psi^{p}}(S)$ denote the subset of probability measures $\mu\in\mathcal{P}(S)$ such that $\mu(\psi^{p}(\cdot,y))<\infty$ for some and hence all given $y\in S$. Next, let $\mu\in\mathcal{P}_{V}(S)\subseteq\mathcal{P}_{\psi^{p}}(S)$, $r>r_{0}\vee r_{\epsilon}$ and $\iota\in[1/2,1[$. Then, by using the first part of Theorem 2.3, there exists a constant $\lambda\in]0,1[$ such that
$D_{\kappa_{\iota,\rho_{\epsilon}(r)}}\Big(\mu P_{n},\mu P_{n-1}\Big)\leq\lambda^{n}D_{\kappa_{\iota,\rho_{\epsilon}(r)}}(\mu P,\mu),\quad\forall n\geq 1.$
Under the assumptions imposed on the semi-distance function $\kappa$ we have
$\ \psi^{p}\leq\Big\|\frac{\psi^{p}}{\kappa}\Big\|~\kappa\leq\Big\|\frac{\psi^{p}}{\kappa}\Big\|\,\Big\|\frac{\kappa}{\varpi_{V}}\Big\|~\varpi_{V}\leq\frac{1}{\rho}\Big\|\frac{\psi^{p}}{\kappa}\Big\|\,\Big\|\frac{\kappa}{\varpi_{V}}\Big\|\varpi_{\rho},\quad\forall\rho>0$
where the last inequality holds by (2.5). This shows that there exists a constant $c>0$ such that $c~\psi^{p}\leq\kappa_{\iota,\rho_{\epsilon}(r)}$, and thus
$\mathbb{W}_{\psi,p}\left(\mu P_{n},\mu P_{n-1}\right)^{p}=D_{\psi^{p}}\left(\mu P_{n},\mu P_{n-1}\right)\leq c^{-1}\lambda^{n}D_{\kappa_{\iota,\rho_{\epsilon}(r)}}(\mu P,\mu),\quad\forall n\geq 1.$ (2.16)
We can now conclude that the mapping $\mu\in\mathcal{P}_{\psi^{p}}(S)\mapsto\mu P$ admits a unique fixed point $\pi\in\mathcal{P}_{\psi^{p}}(S)$ by following the arguments used to prove the Banach fixed point theorem. More precisely, using (2.16) we can easily check that the sequence $(\mu P_{n})_{n\geq 1}$ is Cauchy in $(\mathcal{P}_{\psi^{p}}(S),\mathbb{W}_{\psi,p})$ and thus, since this space is complete (see e.g. Theorem 20.1.8 in [28]), the sequence $(\mu P_{n})_{n\geq 1}$ admits a limit $\pi\in\mathcal{P}_{\psi^{p}}(S)$. The measure $\pi$ is a fixed point of the mapping $\mu\in\mathcal{P}_{\psi^{p}}(S)\mapsto\mu P$ since this mapping is continuous on $(\mathcal{P}_{\psi^{p}}(S),\mathbb{W}_{\psi,p})$. To show that the fixed point $\pi$ is unique assume that there exists another measure $\pi^{\prime}\in\mathcal{P}_{\psi^{p}}(S)$ such that $\pi^{\prime}P=\pi^{\prime}$. In this scenario, for all $n\geq 2$ we have
$D_{\kappa_{\iota,\rho_{\epsilon}(r)}}\left(\pi,\pi^{\prime}\right)=D_{\kappa_{\iota,\rho_{\epsilon}(r)}}\left(\pi P_{n},\pi^{\prime}P_{n-1}\right)=D_{\kappa_{\iota,\rho_{\epsilon}(r)}}\left(\pi P_{n-1},\pi^{\prime}P_{n-1}\right)\leq\lambda D_{\kappa_{\iota,\rho_{\epsilon}(r)}}\left(\pi,\pi^{\prime}\right)$
and thus, since $\lambda<1$, we must have $D_{\kappa_{\iota,\rho_{\epsilon}(r)}}\left(\pi,\pi^{\prime}\right)=0$. Since $\kappa_{\iota,\rho_{\epsilon}(r)}$ is a $V$-semi-distance, it follows that $\pi=\pi^{\prime}$, which concludes to show that the mapping $\mu\in\mathcal{P}_{\psi^{p}}(S)\mapsto\mu P$ has a unique fixed point and thus that $P$ admits a unique invariant distribution $\pi$. This ends the proof of the corollary.
Last but not least, consider Markov integral operators $(K,L)$ and some l.s.c. function $W$, lower bounded away from zero and satisfying (2.1) for some constants $\epsilon\in]0,1[$ and $c\geq 0$. Also assume there exists a map $\alpha:r\in[r_{0},\infty[\mapsto\alpha(r)\in]0,1[$, for some $r_{0}\geq 1$, such that for any $\varpi_{W}(x_{1},x_{2})\leq r$ and $\varpi_{V}(y_{1},y_{2})\leq r$ we have
$\|\delta_{x_{1}}K-\delta_{x_{2}}K\|_{\tiny tv}\vee\|\delta_{y_{1}}L-\delta_{y_{2}}L\|_{\tiny tv}\leq 1-\alpha(r).$ (2.17)
Following the discussion just above Theorem 2.3, we assume without loss of generality that (2.1) and thus (2.17) holds with $W\wedge V\geq 1/2$ and $c=1/2$. In this context, the Markov $V$-operator $P:=KL$ satisfies (H1) and (H2) for the discrete distance $\kappa=\varphi_{0}$ so that the contraction estimate (2.14) is satisfied. However, as we now show, for this specific definition of $P$ a more precise estimate can be obtained. Following word-for-word the proof of (2.14) we check that
$\raisebox{2.0pt}{$\chi$}_{V_{\rho_{\epsilon}(r)},W_{\rho_{\epsilon}(r)}}(K)\vee\raisebox{2.0pt}{$\chi$}_{W_{\rho_{\epsilon}(r)},V_{\rho_{\epsilon}(r)}}(L)\leq 1-\delta_{\epsilon}(r)~\frac{\alpha(r)}{2}$
where the function $W_{\rho}$ is defined as $V_{\rho}$ by replacing $V$ by $W$. Then, by using (2.7), we conclude that
[…]
with $\varphi_{\rho}$ defined in (2.3) and (2.4).
2.5 Exponential decay of $\phi$-Dobrushin coefficients
Theorem 2.3 introduces a class of $V$-semi-distances $\mathcal{F}_{1}$ such that, for any $\phi\in\mathcal{F}_{1}$, there is a constant $\lambda\in]0,1[$ for which $\beta_{\phi}(P_{n})\leq\lambda^{n}$ for all $n\geq 0$. The goal of this subsection is to establish that an exponential decay of $\beta_{\phi}(P_{n})$ holds for a larger class of $V$-semi-distances. Using the comparison results stated in Lemma 2.2, this goal can be easily achieved from the conclusions of Theorem 2.3.
Specifically, we obtain the following result, expressed in terms of the semi-distances $(\varphi_{V},\phi_{V},\kappa_{\iota,V})$ introduced in (1.4) and (1.5).
Theorem 2.5.
Assume that (H1) holds. Then,
  • if (H2) is met for some $V$-semi-distance $\kappa$, for any $\iota\in[1/2,1[$ there exist constants $c\geq 0$ and $\lambda\in]0,1[$ such that
    $\beta_{\kappa_{\iota,V}}(P_{n})\leq c\,\lambda^{n},\quad\forall n\geq 0.$ (2.19)
  • if (H2) is met for the $V$-semi-distance $\kappa=\varphi_{0}$, there exist constants $c\geq 0$ and $\lambda\in]0,1[$ such that
    $\beta_{\varphi_{V}}(P_{n})\leq c\,\lambda^{n},\quad\forall n\geq 0.$ (2.20)
    Moreover, for any $V$-semi-distance $\phi$, there exist constants $c\geq 0$ and $\lambda\in]0,1[$ such that
    $\beta_{\phi_{V}}(P_{n})\leq c\,\lambda^{n},\quad\forall n\geq 0.$ (2.21)
Proof.
Using (2.8) and the remark made just before Theorem 2.3, to prove the theorem we can without loss of generality assume that (H1) holds with $c=1/2$.
To prove (2.19) let $\iota\in[1/2,1]$ and remark first that, using (2.5), for any $\rho\in]0,1[$ we have
$\rho^{1-\iota}~\kappa_{\iota,V}=\rho^{1-\iota}~\varpi_{V}^{1-\iota}\kappa^{\iota}\leq\kappa_{\iota,\rho}=\varpi_{\rho}^{1-\iota}\kappa^{\iota}\leq(1+\rho)^{1-\iota}~\varpi_{V}^{1-\iota}\kappa^{\iota}=(1+\rho)^{1-\iota}~\kappa_{\iota,V},$
and thus, by using the result (2.10) presented in Lemma 2.2,
$\beta_{\kappa_{\iota,V}}(P_{n})\leq(1+1/\rho)^{1-\iota}~\beta_{\kappa_{\iota,\rho}}(P)^{n},\quad\forall n\geq 1,\quad\forall\rho\in]0,1[,\quad\forall r\geq r_{0}.$
Then, (2.19) follows from result (2.13) of Theorem 2.3. Next, by using (2.5) and (2.10), we have
$\beta_{\varphi_{V}}(P_{n})\leq(1+1/\rho)~\beta_{\varphi_{\rho}}(P)^{n},\quad\forall n\geq 1,\quad\forall\rho>0,\quad\forall r\geq r_{0}$
and (2.20) follows from result (2.14) of Theorem 2.3. Finally, to show (2.21) remark that for any constants $\rho>0$ and $\iota\in[0,\rho]$ we have, using (2.5),
$\iota~\big(\varphi_{V}+\phi)\leq\varphi_{\rho}+\iota\phi\leq((1+\rho)\vee\iota)\big(\varphi_{V}+\phi)$
and thus, by (2.10), for any constants $\rho>0$ and $\iota\in[0,\rho]$ we have
$\beta_{\phi+\varphi_{V}}(P_{n})\leq\frac{((1+\rho)\vee\iota)}{(\rho\wedge\iota)}~\beta_{\iota\phi+\varphi_{\rho}}(P)^{n},\quad\forall n\geq 1,\quad\forall r\geq r_{0}.$
Then, (2.21) follows from result (2.15) of Theorem 2.3. This ends the proof of the theorem. ∎
2.6 Exponential decay of the $(\varphi_{V},\phi)$-Dobrushin contraction coefficient
The previous Theorem 2.5 introduces a class of $V$-semi-distances $\mathcal{F}_{2}$ such that, for any $\phi\in\mathcal{F}_{2}$, the $V$-Dobrushin contraction coefficient $\beta_{\phi}(P_{n})$ converges to zero exponentially fast as $n\rightarrow\infty$. By using the contraction estimates provided in this theorem, as well as the trivial observation that for any $V$-semi-distances $\varphi$ and $\phi$ we have
$\varphi\leq\phi\implies D_{\varphi}(\mu_{1},\mu_{2})\leq D_{\phi}(\mu_{1},\mu_{2}),$
we can prove an exponential decay of the $(\varphi_{V},\phi)$-Dobrushin contraction coefficient $\beta_{\varphi_{V},\phi}(P_{n})$ for a larger class of $V$-semi-distances.
Theorem 2.6.
Assume that (H1) is met and that (H2) holds for some $V$-semi-distance $\kappa$. Then, for any $\iota\in]1/2,1]$, there exist constants $c>0$ and $\lambda\in]0,1[$ such that
$\beta_{\varphi_{V},\kappa_{\iota,V}}(P_{n})\leq c~\lambda^{n},\quad\forall n\geq 0.$ (2.22)
with the semi-distance $\kappa_{\iota,V}$ as defined in (1.5). Moreover, if $\beta_{\kappa,\varphi_{0}}(P)<\infty$ then, for any $V$-semi-distance $\phi$, there exists a constant $c^{\prime}\geq 0$ such that
$\beta_{\varphi_{V},\phi}(P_{n})\leq c^{\prime}~\lambda^{n},\quad\forall n\geq 0.$
with the weighted discrete metric $\varphi_{V}$ as defined in (1.4).
Proof.
See Section 4.5. ∎
Remark 2.7.
If $\kappa$ is such that (H2) holds then $\beta_{\kappa}(P)\leq 1$ and we have $\beta_{\kappa,\varphi_{0}}(P)<\infty$ if there exists a constant $c>0$ such that $\|\mu_{1}-\mu_{2}\|_{tv}\leq cD_{\kappa}(\mu_{1},\mu_{2})$ for all $\mu_{1},\mu_{2}\in\mathcal{P}_{V}(S)$.
It is important to remark that, unlike in the first part of Theorem 2.5, in the first part of Theorem 2.6 the case $\iota=1$ is allowed. In this scenario, the first part of Theorem 2.6 states that
[…]
It is also worth noting that the second part of Theorem 2.6 implies that if (H2) holds for the discrete metric (that is with $\kappa=\varphi_{0}$) then the $V$-Dobrushin contraction coefficient $\beta_{\varphi_{V}}(P_{n})$ converges to zero exponentially fast as $n\rightarrow\infty$.
2.6.1 Application 1: Wasserstein exponential contraction estimates
As a first application of Theorem 2.6 we show below how it can be used to establish Wasserstein exponential contraction estimates. To this aim we assume that $(S,\|\cdot\|)$ is a finite dimensional normed vector space and we let $\psi_{S}(x,y)=\|x-y\|$.
Assume first that (H1) holds with $V$ a polynomial-type Lyapunov function, that is with $V(x)=\frac{1}{2}+\|x\|^{p}$ for all $x\in S$ and for some $p>0$. In this case, as proved in Section 4.6, we have the estimate
$\varpi_{V}(x,y)\geq c_{p}~\psi_{S}(x,y)^{p}\quad\mbox{\rm with}\quad c_{p}:=2^{-(p-1)_{+}}$ (2.24)
showing that the function $\psi_{S}^{p}$ is a $V$-semi-distance. In this scenario, Theorem 2.6 ensures that if (H2) is met either with a $V$-semi-distance function $\kappa\geq\psi^{p}_{S}$ or with $\kappa=\varphi_{0}$ then there exist constants $c^{\prime}_{p}>0$ and $\lambda_{p}\in]0,1[$ such that, for all $\mu_{1},\mu_{2}\in\mathcal{P}_{V}(S)$,
$\mathbb{W}_{\psi_{S},p}\left(\mu_{1}P_{n},\mu_{2}P_{n}\right)^{p}=D_{\psi_{S}^{p}}\left(\mu_{1}P_{n},\mu_{2}P_{n}\right)\leq c^{\prime}_{p}~\lambda_{p}^{n}~\|\mu_{1}-\mu_{2}\|_{V},\quad\forall n\geq 0.$
Assume now that (H1) holds with $V$ an exponential-type Lyapunov function, that is with $V(x)=\exp{(\delta\|x\|)}$ for all $x\in S$ and some $\delta>0$. In this case, we show in Section 4.6 that
$\begin{array}[]{l}\displaystyle\varpi_{V}(x,y)\geq\varpi_{e}(x,y)\geq c_{p}\,\psi_{S}(x,y)^{p},\quad\mbox{\rm with}\quad c_{p}:=\frac{\delta^{p}}{2^{p-1}p!},\quad\forall p\geq 1\end{array}$ (2.25)
and where the function $\varpi_{e}$ is defined by
$\varpi_{e}(x,y)=2\left(\exp{\left(\frac{\delta\|x-y\|}{2}\right)}-1\right),\quad x,y\in S.$ (2.26)
By (2.25), the function $\varpi_{e}$ is a $V$-semi-distance, and it follows from Theorem 2.6 that if (H2) holds either for some $V$-semi-distance $\kappa\geq\varpi_{e}$ or with $\kappa=\varphi_{0}$ then there exist constants $c^{\prime}>0$ and $\lambda\in]0,1[$ such that, for all $\mu_{1},\mu_{2}\in\mathcal{P}_{V}(S)$,
$\mathbb{W}_{\psi_{S},p}\left(\mu_{1}P_{n},\mu_{2}P_{n}\right)^{p}\leq c_{p}^{-1}D_{\varpi_{e}}\left(\mu_{1}P_{n},\mu_{2}P_{n}\right)\leq c_{p}^{-1}\,c^{\prime}~\lambda^{n}\|\mu_{1}-\mu_{2}\|_{V},\quad\forall n\geq 0,\quad\forall p\geq 1.$
2.6.2 Application 2: Continuous time Markov semigroups
As a second application of Theorem 2.6 we show below how it can be used to obtain exponential contraction estimates for a continuous time Markov semigroup $(Q_{t})_{t\in[0,\infty[}$, which satisfies the semigroup property $Q_{s+t}=Q_{s}Q_{t}$ for any $s,t\geq 0$.
Let $\cal G$} denote the generator of the semigroup, defined on some domain ${\cal D}({\cal G})$, and assume that there exist a l.s.c. function $V\in{\cal D}({\cal G})$ with $V\geq 1/2$ and some parameters $a_{0},a_{1}>0$ such that
${\cal G}(V)\leq-a_{0}~V+a_{1}.$ (2.27)
In this context, for any $h>0$ we have the following rather well known estimate (see for instance [4])
$Q_{h}(V)\leq\epsilon_{h}V+c_{h}\quad\mbox{\rm with}\quad\epsilon_{h}=(1+a_{0}h)^{-1}<1\quad\mbox{\rm and}\quad c_{h}=a_{1}h.$ (2.28)
We now let $h\in]0,1]$ be fixed and $P=Q_{h}$. Then, by (2.28), condition (H1) is met for the function $V$ such that (2.27) holds. We now assume that, for $P=Q_{h}$, condition (H2) holds for some semi-distance $\kappa$. Noting that $(P_{n})_{n\geq 0}=(Q_{nh})_{n\geq 0}$, that is that $(Q_{nh})_{n\geq 0}$ is the discrete generation semigroup of Markov $V$-operators associated with the Markov $V$-operator $P=Q_{h}$, it follows from Theorem 2.6 that there exist constants $c_{h}>0$ and $\lambda_{h}\in]0,1[$ such that
$\beta_{\varphi_{V},\kappa}(Q_{nh})\leq c_{h}\lambda_{h}^{n},\quad\forall n\geq 0.$ (2.29)
To extend the contraction estimate (2.29) obtained for the discrete time Markov semigroup $(Q_{nh})_{n\geq 1}$ to the continuous time Markov semigroup $(Q_{t})_{t\in[0,\infty[}$ remark first that
$Q_{t}=Q_{h\{t/h\}}Q_{h\lfloor t/h\rfloor}=Q_{h\{t/h\}}P_{\lfloor t/h\rfloor},\quad\forall t>0$ (2.30)
where $\lfloor s\rfloor$ and $\{s\}:=s-\lfloor s\rfloor<1$ stands for the integer part and the fractional part of $s>0$. In addition, remark that since for any $t>0$ we have $\{t/h\}h<h\leq 1$, it follows from (2.28) that
$\|Q_{\{t/h\}h}(V)/V\|\leq\sup_{0\leq h\leq 1}\|Q_{h}(V)/V\|<\infty,\quad\forall t>0$
implying that there exists a constant $\iota>0$ for which we have
$\|\mu_{1}Q_{h\{t/h\}}-\mu_{2}Q_{h\{t/h\}}\|_{V}\leq\iota~\|\mu_{1}-\mu_{2}\|_{V}\quad\forall t>0,\quad\forall\mu_{1},\mu_{2}\in\mathcal{P}_{V}(S).$ (2.31)
By combining (2.29)-(2.31), we obtain the following contraction estimate for the continuous time Markov semigroup $(Q_{t})_{t\in[0,\infty[}$
$\beta_{\varphi_{V},\kappa}(Q_{t})\leq\iota\,c_{h}~\lambda_{h}^{\lfloor t/h\rfloor},\quad\forall t>0.$ (2.32)
Equivalently, for any $t>0$ we have the exponential contraction estimate
$c_{\iota,h}:=\iota\,c_{h}~\lambda_{h}^{-1}\Longrightarrow\beta_{\varphi_{V},\kappa}(Q_{t})\leq c_{\iota,h}~e^{-\varsigma_{h}t}\quad\mbox{\rm with}\quad\varsigma_{h}:=-\frac{1}{h}\log{\lambda_{h}}>0.$ (2.33)
Moreover, if $\beta_{\kappa,\varphi_{0}}(Q_{h})<\infty$ then, by using the second part of Theorem 2.6 as well as (2.30)-(2.31), we readily obtain that for any $V$-semi-distance $\phi$ there exist a constant $c^{\prime}_{\iota,h}\geq 0$ such that, with $\varsigma_{h}>0$ as in (2.33),
$\beta_{\varphi_{V},\phi}(Q_{t})\leq c^{\prime}_{\iota,h}~e^{-\varsigma_{h}t},\quad\forall t>0.$ (2.34)
2.7 Discussion
Local Lyapunov techniques w.r.t. weighted total variation norms go back to the beginning of the 1990s with the pioneering articles by Meyn and Tweedie [56, 57, 58]. They were further developed in the articles [4, 21, 40, 42, 67] and in the book [22], simplifying the Foster-Lyapunov methodologies and the small-sets return times estimation techniques presented in [34, 56, 57, 58, 59]. In this context, the present article extends the operator theoretic approach developed in [4, 22] for weighted total variation norms to a larger class of Kantorovitch semi-distances.
As underlined in [31], in practical applications it is usually unclear how to choose the semi-distance function $\phi$ to ensure that $\beta_{\phi}(P)<1$. To answer this question, more or less involved and carefully designed semi-distance functions combined with appropriate couplings depending on the model at hand have been proposed in the literature, see for instance the refined analysis developed in the series of articles [8, 30, 31, 43, 46]. It is interesting to remark that the semi-distance function in the strict contraction estimate (2.13) provided in Theorem 2.3 is similar to the one in Theorem 2.6 of the article [8] (dedicated to Hamiltonian Monte Carlo) and to the one in Theorem 4.8 of [43]. On the other hand, the semi-distance in the strict contraction estimate (2.15) given in Theorem 2.3 is similar to the one in Theorem 2.3 of [31] and in Theorem 2.3 of [46].
The contraction estimate (2.14) ensures the existence of a single invariant probability measure $\mu=\mu P$, while the second part of Theorem 2.6 establishes the exponential contraction estimate $\|\mu_{1}P_{n}-\mu_{2}P_{n}\|_{V}\leq c\lambda^{n}\|\mu_{1}-\mu_{2}\|_{V}$. It is well-known that these two properties of $P$ hold under the drift condition (H1) and when the local contraction condition (H2) is met for the discrete metric. However, as reminded with the present paper, the weighted-norm contraction coefficients technique developed in [4, 22] provides a very direct and short proof of these two results.
Corollary 2.4 shows that the contraction estimate (2.13) also ensures the existence of a unique invariant distribution for $P$ under mild assumptions on the semi-distance $\kappa$ for which the local contraction condition (H2) holds. In this scenario, the result of Corollary 2.4 is almost identical to part (i) of Theorem 40.4.5 in [28] (see also Proposition 2.9 in [63]), but the operator theoretic approach used in our article provides a very direct and concise proof.
Under assumptions which are essentially equivalent to assuming that (H1) holds and that (H2) is met for some semi-distance $\kappa$, an exponential convergence result of the form $D_{\kappa}(\mu_{1}P_{n},\mu_{2}P_{n})\leq c\,\lambda^{n}\big(\mu_{1}(V)+\mu_{2}(V)\big)$ is obtained for instance in [63] (see in particular Corollary 2.1 and Theorem 2.6) and in Theorem 40.4.5 of the book [28]. Theorem 2.6 improves upon this result by establishing an exponential contraction estimate of the form $D_{\kappa}(\mu_{1}P_{n},\mu_{2}P_{n})\leq c\,\lambda^{n}\|\mu_{1}-\mu_{2}\|_{V}$.
A contraction result for a continuous time Markov semigroup $(Q_{t})_{t\in]0,\infty[}$ is obtained in [43], Theorem 4.8. Assuming the existence of a Lyapunov function $V$, this theorem shows that, for a symmetric and semi-distance $\kappa$ bounded by one, and with $\iota=1/2$ and $\rho=1$, the inequality $\beta_{\kappa_{\iota,\rho}}(Q_{t})<1$ holds for $t$ sufficiently large. However, this result further assumes that there exists a $t^{\prime}>0$ such that $\sup_{x\neq y}\ell_{\kappa}(Q_{t^{\prime}})(x,y)<1$ which, as proved in Lemma 1.1, is equivalent to assuming that $\beta_{\kappa}(Q_{t^{\prime}})<1$. In the context of $\mathbb{R}^{n}$-valued diffusions, an exponential contraction estimate is obtained in [30], assuming the existence of a Lyapunov function $V$ satisfying the geometric drift condition (2.27) and some additional key growth conditions (see Assumptions 2.3 in [30]). Specifically, Theorem 2.1 in [30] shows that $\beta_{\phi}(Q_{t})\leq\lambda^{t}$ for some constant $\lambda\in]0,1[$, where $\phi(x,y)=\varpi(\|x-y\|)+\varphi_{\rho}(x,y)$ for some judiciously chosen bounded non-decreasing concave continuous functions $\varpi$ and constant $\rho>0$. The proof of this result combines diffusion reflection coupling techniques with a general Ito-Tanaka type formula and martingales local times techniques to estimate the time spent by the processes on different level sets. By contrast, for the same diffusion model as the one considered in [30] and without any additional assumptions, exponential contraction estimates can be easily deduced from Theorem 2.6, following the discussion in Section 2.6.2 (see Section 3.5 for details). We however stress that these estimates do not establish that each Markov operator $Q_{t}$ is a contraction, unlike the result obtained in [30] (again, see Section 3.5 for details).
3 Applications
In what follows we let ${\cal B}_{0}(S)=\{{\cal V}:1/{\cal V}\in B_{\infty}(S)\}$ be the sub-algebra of (bounded) positive functions $\Theta$, which are locally lower bounded and vanish at infinity in the sense that, for any $0<\epsilon<\|\Theta\|$, the set $C_{\epsilon}:=\{\Theta\geq\epsilon\}$ is a non-empty compact set. We recall that if a function $W>0$ satisfies
$P(W)/W\leq\Theta\text{ for some $\Theta\in{\cal B}_{0}(S)$}$ (3.1)
then, for any $0<\epsilon<\|\Theta\|\wedge 1$, we have
$P(W)\leq\epsilon~1_{S-C_{\epsilon}}~W+1_{C_{\epsilon}}~\Theta~W\leq\epsilon~W+c_{\epsilon}\quad\mbox{\rm with}\quad c_{\epsilon}:=\|\Theta\|~\sup_{x\in C_{\epsilon}}W(x).$ (3.2)
Consequently, if (3.1) holds for a l.s.c. function $W$ then (H1) holds with $V=1+W$.
3.1 Some illustrative examples
3.1.1 Markov chains on the unit open interval with bounded densities
We let $S=]0,1[$, $\psi_{e}(x,y)=|x-y|$ be the Euclidean metric and, for any $\iota\in]0,1]$ we let
$\psi_{\iota}(x,y):=\left|\frac{1}{x}-\frac{1}{y}\right|^{\iota}+\left|\frac{1}{1-x}-\frac{1}{1-y}\right|^{\iota}.$ (3.3)
We note that $(S,\psi_{\iota})$ is a complete and separable metric space for any $\iota\in]0,1]$ while the space $(S,\psi_{e})$ is separable but not complete.
Next, given a probability measure $\nu(dy)=g(y)dy\in\mathcal{P}(S)$ having a bounded density $g$ and a Markov transition $Q(x,dy)=q(x,y)dy$ on ]0,1[]0,1[ such that $\|q\|<\infty$, we let
$\begin{array}[]{l}P(x,dy)=x\,Q(x,dy)+(1-x)\nu(dy)\end{array}$ (3.4)
and, for some $\iota\in]0,1[$, we define the continuous function
$W_{\iota}(x)=\frac{1}{x^{\iota}}+\frac{1}{(1-x)^{\iota}}.$ (3.5)
Noting that $W_{\iota}\in{\cal B}_{\infty}(S)$ while $\|P(W_{\iota})\|\leq c:=2(\|q\|+\|g\|)/(1-\iota)<\infty$, it follows that (3.1) holds with $\Theta=1/W_{\iota}$ and thus, by (3.2), condition (H1) holds with $V=W_{\iota}$ (observe that $W_{\iota}$ is bounded below away from zero).
Assume first that the two probability distributions $\nu(dy)$ and $Q(x,dy)$ have disjoint supports for all $x\in S$. In this scenario, we have $\|\delta_{x}P-\delta_{0}P\|_{\tiny tv}=x$ implying that $\beta_{\varphi_{0}}(P)=1$ and therefore (H2) does not hold with the discrete metric, that is with $\kappa=\varphi_{0}$.
Assume now that, for all $x\in S$, the two probability distributions $\nu(dy)$ and $Q(x,dy)$ have $S$ as support, and assume also that their densities are continuous and strictly positive. In this scenario, (H2) holds with $\kappa=\varphi_{0}$ by Lemma 2.1, and Theorem 2.6 therefore ensures that the Kantorovitch semi-distance $D_{\phi}(\mu_{1}P_{n},\mu_{2}P_{n})$ vanishes exponentially quickly as $n\rightarrow\infty$ for any $V$-semi-distance $\phi$. In particular, noting that the complete metric $\psi_{S}=\psi_{\iota}$ is a $V$-semi-distance function, Theorem 2.6 gives the following exponential contraction estimate for the Wasserstein 1-distance:
$\mathbb{W}_{\psi_{S}}\left(\mu_{1}P_{n},\mu_{2}P_{n}\right)\leq c\lambda^{n}\|\mu_{1}-\mu_{2}\|_{V},\quad\forall n\geq 0,\quad\forall\mu_{1},\mu_{2}\in\mathcal{P}_{V}(S).$ (3.6)
Lastly, following an example proposed in [24], Section 2.1, we assume that
$P(x,dy):=\frac{1}{2x}~1_{]0,x]}(y)\,dy+\frac{1}{2(1-x)}~1_{]x,1[}(y)\,dy$
so that the Markov transition $P$ has an unbounded density. In this context, we let $\iota\in]0,1/2[$ and show in Section 4.7 that there exist some parameters $\epsilon\in[0,1[$ and $c\geq 0$ such that
$P(W_{\iota})(x)\leq\epsilon\,W_{\iota}(x)+c$ (3.7)
and thus such that condition (H1) holds with $V=W_{\iota}$. Since $V\in{\cal B}_{\infty}(S)$ while for all $x\in S$ we have $P(x,dy)\geq\frac{1}{2}~1_{]0,1[}(y)~dy$, the local minorization condition (2.2) is satisfied on the compact sub-level sets of $V$ and thus (H2) is met with the $V$-semi-distance $\kappa=\varphi_{0}$. Therefore, noting that the complete metric $\psi_{S}=\psi_{\iota}$ is a $V$-semi-distance, it follows that the exponential contraction estimate (3.6) for the Wassertein 1-distance holds.
3.1.2 A Markov chain on the half-line
Let $S=]0,\infty[$ and $P(x,dy)$ be the Markov transition on the half-line $S=]0,\infty[$ given by
$P(x,dy):=\frac{\delta}{2}~\frac{1}{x}~1_{]0,x]}(y)~dy+\frac{\delta}{2}~g(y-x)~dy+(1-\delta)\nu_{\gamma}(dy).$
In the above display, $g$ and $\nu_{\gamma}$ stand for the Gaussian density and the Weibull distribution associated with a shape parameter $\gamma>0$ defined by
$g(z)=1_{]0,\infty[}(z)~\sqrt{\frac{2}{\pi}}~e^{-z^{2}/2}\quad\mbox{\rm and}\quad\nu_{\gamma}(dz):=1_{]0,\infty[}(z)~(1+\gamma)~z^{\gamma}~e^{-z^{1+\gamma}}~dz.$
We remark that, as in the previous example, the density of $P$ is unbounded and the space $S$ equipped with the Euclidean metric $\psi_{e}(x,y)=|x-y|$ is separable but not complete. It can however be shown that $(S,\psi_{S})$ is a complete and separable metric space, where the metric $\psi_{S}$ is defined, for some $\iota\in]0,1[$, by
$\psi_{S}(x,y):=\left|\frac{1}{x}-\frac{1}{y}\right|^{\iota}+\left|x-y\right|.$ (3.8)
We now let $W(x)=x^{-\iota}+x$ with $\iota$ as in (3.8) and remark that $W\in{\cal B}_{\infty}(S)$. In Section 4.7 we show that there exist some parameters $\epsilon\in]0,1[$ and $c\geq 0$ such that
$P(W)\leq\epsilon\,W+c$ (3.9)
and therefore condition (H1) holds with $V=W$ (observe that $W$ is bounded below away from zero). In addition, since $V\in{\cal B}_{\infty}(S)$ while for all $x\in S$ we have $P(x,dy)\geq(1-\delta)\nu_{\gamma}(dy)$, the local minorization condition (2.2) is satisfied on the compact sub-level sets of $V$ and thus (H2) is met with the $V$-semi-distance function $\kappa=\varphi_{0}$. Therefore, noting that the complete metric $\psi_{S}$ defined in (3.8) is $V$-semi-distance, it follows from Theorem 2.6 that the exponential contraction estimate (3.6) for the Wassertein 1-distance holds.
3.2 Markov chains on domains with a boundary
In this subsection we consider a Markov transition $P(x,dy)=p(x,y)dy$ with a uniformly bounded, strictly positive and continuous density $p(x,y)$ on a open and connected set $S\subset\mathbb{R}^{n}$ having a non-empty boundary $\partial S$. We define the continuous function
$d_{n}(x,\partial S):=\inf{\left\{\|x-y\|~:~y\in\partial S\right\}}$
and, for some $\iota\in]0,1]$, we equip $S$ with the metric
$\psi_{S}(x,y):=\left|\frac{1}{d_{n}(x,\partial S)}-\frac{1}{d_{n}(y,\partial S)}\right|^{\iota}+\|x-y\|^{\iota}.$ (3.10)
It can be shown that $(S,\psi_{S})$ is a complete and separable metric space.
3.2.1 Bounded domains with a locally Lipschitz boundary
Assume that $S$ is bounded and has a locally Lipschitz boundary $\partial S$, and let $W(x)=d_{n}(x,\partial S)^{-\iota}+\|x\|$. In this scenario, we clearly have $W\in{\cal B}_{\infty}(S)$ and, since the function $W$ is integrable over the set $S$ (c.f. Lemma 5.2 in [4]), under the above assumptions on $P$ we have $\|P(W)\|<\infty$. Consequently, (3.1) is met with $\Theta=1/W$ and it follows from (3.2) that (H1) holds with $V=1+W$. On the other hand, since $P$ is assumed to have a strictly positive and continuous density, condition (H2) holds with $\kappa=\varphi_{0}$ by Lemma 2.1. By Theorem 2.6, it follows that for any $V$-semi-distance $\phi$ the Kantorovitch semi-distance $D_{\phi}(\mu_{1}P_{n},\mu_{2}P_{n})$ converges to zero exponentially quickly as $n\rightarrow\infty$. In particular, one can check that the metric $\psi_{S}$ defined in (3.10) is the $V$-semi-distance function, and thus Theorem 2.6 gives the following exponential contraction estimate for the Wasserstein 1-distance:
$\mathbb{W}_{\psi_{S}}\left(\mu_{1}P_{n},\mu_{2}P_{n}\right)\leq c\lambda^{n}\|\mu_{1}-\mu_{2}\|_{V},\quad\forall n\geq 0,\quad\forall\mu_{1},\mu_{2}\in\mathcal{P}_{V}(S).$ (3.11)
3.2.2 Domains with a locally Lipschitz bounded boundary
Assume that $S$ is not necessarily bounded but has a locally Lipschitz bounded boundary $\partial S$. Assume also that there exists a function $W_{1}\in{\cal B}_{\infty}(\mathbb{R}^{n})$ such that $\sup_{x\in S}\|x\|^{\iota}/W_{1}(x)<\infty$ and such that $P(W_{1})\leq\epsilon_{1}W_{1}+c_{1}$ for some parameters $\epsilon_{1}\in[0,1[$ and $c_{1}>0$. In this scenario, condition (H1) holds with $V=1+W_{1}$ but, for this definition of $V$, the metric $\psi_{S}$ defined in (3.10) may not be a $V$-semi-distance. To solve this problem we let $W_{2}(x)=d_{n}(x,\partial S)^{-\iota}\in{\cal B}_{\infty}(S)$ and $c_{2}=\|P(W_{2})\|\leq 1+\|P(W_{2}1_{W_{2}\geq 1})\|$. Using the fact that $\{x\in S:W_{2}(x)\geq 1\}=\{x\in S:d_{n}(x,\partial S)\leq 1\}$ with $\partial S$ a bounded set, and recalling that the density $p(x,y)$ is assumed to be uniformly bounded, by using a similar argument as in the proof of Lemma 5.2 in [4] we can check that $\|P(W_{2}1_{W_{2}\geq 1})\|<\infty$. This implies that $c_{2}<\infty$ and thus, since
$\frac{P(W_{1}+W_{2})}{W_{1}+W_{2}}\leq\frac{\epsilon_{1}W_{1}+(c_{1}+c_{2})}{W_{1}+W_{2}}\leq\epsilon_{1}+\frac{(c_{1}+c_{2})}{W_{1}+W_{2}},$
it follows that (H1) also holds with $V=1+W_{1}+W_{2}\in{\cal B}_{\infty}(S)$. For this definition of $V$ the metric $\psi_{S}$ defined in (3.10) is a $V$-semi-distance and, as in Section 3.2.1, we can apply Lemma 2.1 to conclude that (H2) holds with $\kappa=\varphi_{0}$. Therefore, by Theorem 2.6, the exponential Wasserstein contraction estimate (3.11) holds.
3.3 Block-Gibbs samplers
3.3.1 Conjugate transformations
We consider a locally compact Polish space $S$ endowed with some locally bounded positive measure $\nu$, and we let $(\nu_{g},\nu_{h})$ be Boltzmann-Gibbs probability distributions on $S$ defined by
$\nu_{g}(dx):=\exp{(-g(x))}~\nu(dx)\quad\mbox{\rm and}\quad\nu_{h}(dy):=\exp{(-h(y))}~\nu(dy)$ (3.12)
for some normalized potential functions $g:x\in S\mapsto g(x)\in\mathbb{R}$ and $h:y\in S\mapsto h(y)\in\mathbb{R}$. In addition, we let $M(y,dx)$ be a Markov transition of the form
$M(y,dx)=\exp{(-m(y,x))}~\nu(dx)$ (3.13)
for some normalized potential transition $m:(y,x)\in S^{2}\mapsto m(y,x)\in\mathbb{R}$.
In this context, we have the duality formula
$\pi_{h}(d(x,y)):=\nu_{h}(dy)~M(y,dx)=(\nu_{h}M)(dx)~M^{\star}_{\nu_{h}}(x,dy)$
where the conjugate (a.k.a. adjoint or backward) transition $M^{\star}_{\nu_{h}}$ is given by the Markov transition
$M^{\star}_{\nu_{h}}(x,dy)=K(x,dy):=\frac{d\nu}{d\nu_{h}M}(x)~\nu_{h}(dy)~\exp{(-m(y,x))}\Longrightarrow\nu_{h}MK=\nu_{h}.$ (3.14)
To go one step further in our discussion we let $\pi_{g}$ be the probability measure on $S^{2}$ defined by
$\pi_{g}(d(x,y)):=\nu_{g}(dx)~K(x,dy)=(\nu_{g}K)(dy)~K^{\star}_{\nu_{g}}(y,dx)$
where the conjugate transition $K^{\star}_{\nu_{g}}$ is given by the Markov transition
$K^{\star}_{\nu_{g}}(y,dx)=L(y,dx):=\frac{1}{M\left(\frac{d\nu_{g}}{d\nu_{h}M}\right)(y)}~M(y,dx)~\frac{d\nu_{g}}{d\nu_{h}K}(x)\Longrightarrow\nu_{g}KL=\nu_{g}.$ (3.15)
We stress that large class of forward-backward-type samplers, including proximal samplers [9, 14, 15, 48, 49, 52, 61], Glauber dynamics [13, 38], Sinkhorn Gibbs-loop processes [2, 3, 20] and two block-Gibbs samplers [17, 35, 36, 70] with a prescribed target measure, have a two stages transition $P=KL$ of the product form of a Markov transition $K$ and its adjoint $L$; see also  [1, 12, 37, 39, 64, 69] and references therein.
The stability analysis of these Markov chains are traditionally based on spectral gaps or relative entropies and log-Sobolev inequalities [5, 11, 19, 25, 26, 47], see also the recent article [41] and references therein. The stability of Gibbs samplers in particular classes of statistical conjugate models has also been studied using Lyapunov and Wasserstein approaches, respectively in the articles  [10, 44, 65, 66] and [45]. As underlined in [10] the design of appropriate Lyapunov functions depends on some understanding of the dynamics of the Gibbs sampler which is often implicitly defined. Next section provides a simple method for designing Lyapunov functions for transitions $(K,L)$ of the form (3.14) and (3.15).
3.3.2 Lyapunov functions
Let us assume that the functions $(g,h,m)$ are locally bounded and bounded below by real numbers $(g^{-},h^{-},m^{-})$ and the that two functions $(g,h)$ have compact sub-level sets.
Consider the functions
$g_{\delta}:=\delta g-m_{h}\quad\mbox{\rm and}\quad h_{\delta}:=\delta h-m_{g}$ (3.16)
indexed by $\delta\in]0,1[$ with the integrated functions $(m_{g},m_{h})$ defined by
$m_{h}(x):=\int\nu_{h}(dz)~m(z,x)\quad\text{and}\quad m_{g}(y):=\int\nu_{g}(dz)~m(y,z).$ (3.17)
Note that $m_{h}\wedge m_{g}\geq m^{-}$. We further assume there is some $\delta\in]0,1/2[$ such that the functions $(g_{\delta},h_{\delta})$ are locally bounded with compact sub-level sets and such that
$\nu_{(1-2\delta)g}\left(1\right)\vee\nu_{(1-\delta)h}\left(1\right)<\infty.$ (3.18)
Note that $(g_{\delta},h_{\delta})$ are bounded below by real numbers $(g^{-}_{\delta},h^{-}_{\delta})$. Various examples satisfying the above regularity conditions are discussed in [2].
Example 3.1.
Assume that $(g,h)$ and $(m_{g},m_{h})$ are continuous functions. In addition, assume that there exist some parameters $a_{1},b_{1}\in\mathbb{R}$, $a_{2},b_{2}>0$ and $q\geq p>0$ such that
$m(x,y)\leq a_{1}+a_{2}~(\|x\|^{p}+\|y\|^{p})\quad\mbox{and}\quad g(x)\wedge h(x)\geq b_{1}+b_{2}~\|x\|^{q}.$
In this case (3.18) holds and $(g_{\delta},h_{\delta})$ are locally bounded with compact sub-level sets for any $\delta\in]0,1/2[$ as soon as $q>p$, and for some $\delta\in]0,1/2[$ as soon as $b_{2}/2>a_{2}$ and $q=p$.
Lemma 3.2.
The Markov transitions $(K,L)$ defined in (3.14) and (3.15) satisfy
$\exp{(-\delta g)}~K\left(\exp{(\delta h)}\right)$ $\leq$ $c_{\delta,h}~\exp{\left(-g_{\delta}\right)}$ (3.19)
$\exp{(-\delta h)}~L\left(\exp{(\delta g)}\right)$ $\leq$ $c_{\delta,g}~\exp{\left(-h_{\delta}\right)}$ (3.20)
with the parameters
$c_{\delta,h}:=\exp{(-m^{-})}~\nu_{(1-\delta)h}(1)\quad\mbox{and}\quad c_{\delta,g}:=\nu_{(1-2\delta)g}(1)~\exp{(-g_{\delta}^{-}-\nu_{g}(g))}.$
Proof.
Using Jensen’s inequality we check that
$\frac{d\nu_{h}M}{d\nu}(x)=\int\nu_{h}(dz)\exp{(-m(z,x))}\geq\exp{\left(-m_{h}(x)\right)}$
yielding the estimate
$\exp{(-\delta g(x))}~K(x,dy)~\exp{(\delta h(y))}\leq c_{0}~\exp{\left(-g_{\delta}(x)\right)}~\nu_{(1-\delta)h}(dy)$ (3.21)
with $c_{0}:=\exp{(-m^{-})}$. We end the proof of (3.19) integrating both sides of (3.21).
To check (3.20), note that
$M\left(\frac{d\nu_{g}}{d\nu_{h}M}\right)(y)=\int~\nu_{g}(dx)~\exp{\left(g(x)-m(y,x)+\log{\frac{d\nu_{g}}{d\nu_{h}M}(x)}\right)}.$
Then, using Jensen’s inequality, we check that
$M\left(\frac{d\nu_{g}}{d\nu_{h}M}\right)(y)\geq\exp{\left(\nu_{g}(g)-m_{g}(y)+\nu_{g}\left(\log{\frac{d\nu_{g}}{d\nu_{h}M}}\right)\right)}\geq\exp{\left(\nu_{g}(g)-m_{g}(y)\right)}.$
This yields the estimate
$L(y,dx)\leq\exp{\left(-\nu_{g}(g)+m_{g}(y)\right)}~e^{-m(y,x)}~\nu_{g}(dx)~\frac{d\nu}{d\nu_{h}M}(x)$
from which we check that
$e^{-\delta h(y)}L(y,dx)e^{\delta g(x)}\leq c_{1}~\exp{\left(-h_{\delta}(y)\right)}~\nu_{(1-2\delta)g}(dx)$ (3.22)
with $c_{1}:=\exp{(-g_{\delta}^{-}-\nu_{g}(g))}$. We end the proof of (3.20) integrating both sides of (3.22). ∎
Lemma 3.2 ensures that transitions $(K,L)$ defined in (3.14) and (3.15) satisfy the drift conditions in (2.1) with $(W,V)=(\exp{(\delta h)},\exp{(\delta g)})$. To check this claim, given some $\epsilon\in]0,1[$, consider some $r>0$ such that
$(c_{\delta,h}\vee c_{\delta,g})\exp{\left(-r\right)}\leq\epsilon.$
Then, by (3.19) and (3.20), the drift conditions in (2.1) are satisfied with $c=c_{1,\delta}(r)\vee c_{2,\delta}(r)$ and the parameters
$c_{1,\delta}(r):=c_{\delta,h}~\exp{\left(\sup_{g_{\delta}(x)\leq r}|\delta g(x)-g_{\delta}(x)|\right)}~\mbox{\rm and}~c_{2,\delta}(r):=c_{\delta,g}~\exp{\left(\sup_{h_{\delta}(x)\leq r}|\delta h(x)-h_{\delta}(x)|\right)}.$
3.3.3 Local minorizations
For any $r>1$, consider the level sets
$C_{V}(r):=\{V\leq r\}\quad\mbox{\rm and}\quad C_{W}(r):=\{W\leq r\}.$
Since $S$ can be exhausted respectively by the compact sub-levels sets $C_{V}(r)$ and $C_{W}(r)$, there exists some $r_{0}>1$ such that $\nu(C_{V}(r))\wedge\nu(C_{W}(r))>0$ for $r=r_{0}$ and thus for any $r\geq r_{0}$. This ensures that there exists some $r_{0}>0$ such that $\nu_{h}(C_{W}(r))\wedge\nu_{g}(C_{V}(r))>0$ for any $r\geq r_{0}$, and we set
$\nu_{g,r}(dx):=\frac{\nu_{g}(dx)~1_{C_{V}(r)}(x)}{\nu_{g}(C_{V}(r))}\quad\mbox{\rm and}\quad\nu_{h,r}(dy):=\frac{\nu_{h}(dy)~1_{C_{W}(r)}(y)}{\nu_{h}(C_{W}(r))},\quad\forall r\geq r_{0}.$
Lemma 3.3.
For any $r\geq r_{0}$ there exists some $\alpha(r)\in]0,1[$ such that for any $x\in C_{V}(r)$ and $y\in C_{W}(r)$ we have
$K(x,dy)~\geq\alpha(r)~\nu_{h,r}(dy)\quad\mbox{and}\quad L(y,dx)\geq~\alpha(r)~\nu_{g,r}(dx).$ (3.23)
Proof.
Observe that, for any $x\in C_{V}(r)$, we have
$\begin{array}[]{l}\displaystyle~K(x,dy)\geq\exp{(m^{-})}~\nu_{h}(dy)~\exp{(-m(y,x))}~1_{C_{W}(r)}(y)\geq\exp{(a(r))}~\nu_{h}(dy)~1_{C_{W}(r)}(y)\end{array}$
with the parameter
$a(r):=m^{-}-\sup_{(x,y)\in(C_{V}(r)\times C_{W}(r))}m(y,x),$
which implies that
$K(x,dy)~\geq\alpha_{h}(r)~\nu_{h,r}(dy)\quad\mbox{\rm with}\quad\alpha_{h}(r):=\exp{(a(r))}~\nu_{h}(C_{W}(r)),\quad\forall x\in C_{V}(r).$
In the same vein, we have
$M(y,dx)~\frac{d\nu_{g}}{d\nu_{h}M}(x)=\frac{\nu_{g}(dx)~e^{-m(y,x)}}{\int\nu_{h}(dy)~e^{-m(y,x)}}\geq e^{m^{-}}~\nu_{g}(dx)~e^{-m(y,x)}$
and thus, by Jensen’s inequality,
$\begin{array}[]{rcl}\displaystyle~M(y,dx)~\frac{d\nu_{g}}{d\nu_{h}M}(x)&\leq&e^{m^{-}}~\nu_{g}(dx)~e^{m_{h}(x)}\\ &=&e^{m^{-}}~\nu_{(1-\delta)g}(dx)~e^{-g_{\delta}(x)}\leq\exp{(m^{-}-g_{\delta}^{-})}~\nu_{(1-\delta)g}(dx).\end{array}$
For any $y\in C_{W}(r)$, this implies that
$\begin{array}[]{l}\displaystyle L(y,dx)\geq~\alpha_{g}(r)~\nu_{g,r}(dx)\end{array}$
with the parameters
$\alpha_{g}(r):=\nu_{g}(C_{V}(r))~\frac{\exp{(b(r))}}{\nu_{(1-\delta)g}(1)}\quad\mbox{\rm and}\quad b(r):=g^{-}_{\delta}-\sup_{(x,y)\in(C_{V}(r)\times C_{W}(r))}m(y,x).$
We conclude that the two inequalities in (3.23) are satisfied with $\alpha(r):=\alpha_{g}(r)\wedge\alpha_{h}(r)$. This ends the proof of the lemma. ∎
The local minorization estimates (3.23) ensures that for any $r\geq r_{0}$, $(x_{1},x_{2})\in C_{V}(r)^{2}$ and $(y_{1},y_{2})\in C_{W}(r)^{2}$ we have
$\|\delta_{x_{1}}K-\delta_{x_{2}}K\|_{\tiny tv}\vee\|\delta_{y_{1}}L-\delta_{y_{2}}L\|_{\tiny tv}\leq 1-\alpha(r)$
from which we conclude that condition (2.17) is met. Recalling that (2.1) is also met, it follows that the contraction estimate (2.18) is holds with $P=KL$, where for any $\rho>0$ the semi-distance $\varphi_{\rho}$ is defined in (2.3) and (2.4).
3.4 Iterated random functions
In this subsection we let $\cal Z$} be a measurable space equipped with some probability distribution $\mu$, $(S,\|\cdot\|)$ be a normed vector space, and we assume that
$P(x,dy)=\mathbb{P}\left(F_{Z}(x)\in dy\right),\quad Z\sim\mu$ (3.24)
for some function $F:(x,z)\in(S\times{\cal Z})\mapsto F_{z}(x)\in S$. In this scenario the Markov semigroup $(P_{n})_{n\geq 0}$ takes the form
$P_{n}(x,dy)=\mathbb{P}\left(X_{n}(x)\in dy\right)\quad\mbox{\rm with}\quad X_{n}(x)=F_{Z_{n}}(X_{n-1}(x))=\left(F_{Z_{n}}\circ\ldots\circ F_{Z_{1}}\right)(x)$
where the $Z_{n}$’s are independent copies of $Z$. Markov chains are often constructed by iterating random functions, and in this framework it is known that the semigroup is stable as soon as the random functions exhibit a certain average contraction, see for instance [24, 58, 68].
3.4.1 Some general local contraction results
Assume that the function $x\mapsto F_{z}(x)$ is smooth for any $z\in{\cal Z}$ and denote by $\nabla_{1}F_{z}(x)$ its gradient, and assume that the function $\|x-y\|^{p}$ is a $V$-semi-distance for some $p>0$. In this context conditions (H2) is met with the semi-distance $\kappa(x,y):=\|x-y\|^{p}$ as soon as, for all $r$ large enough, there exists some $\alpha_{p}(r)\in]0,1[$ such that
$\varpi_{V}(x,y)\leq r\Longrightarrow\mathbb{E}\left(\|F_{Z}(x)-F_{Z}(y)\|^{p}\right)\leq(1-\alpha_{p}(r))\|x-y\|^{p}.$ (3.25)
Notice that (3.25) holds with $p=1$ if $\sup_{x\in\mbox{\rm co}\left(\{V\leq r\}\right)}\mathbb{E}\left(\|\nabla_{1}F_{Z}(x)\|\right)<1$, where $\mbox{\rm co}\left(\{V\leq r\}\right)$ stands for the convex hull of the compact sub-level sets $\{V\leq r\}$, assuming that $V\in{\cal B}_{\infty}(S)$. It is also worth noting that if $F_{z}=F+z$ for some function $F:x\in S\mapsto F(x)\in S$ then condition (3.25) depends only on $F$. In this case, (H2) is met with $\kappa(x,y):=\|x-y\|^{p}$ for any distribution $\mu$ of $Z$.
Alternatively, if $V\in{\cal B}_{\infty}(S)$ and $P(x,dy)=p(x,y)\,\nu(dy)$ for some continuous and strictly positive transition density $p(x,y)$ then, by Lemma 2.1, (H2) holds for the semi-distance $\kappa=\varphi_{0}$.
3.4.2 A general approach to construct Lyapunov functions
We now assume that $F_{Z}=F+G_{Z}$ for some functions $F:x\in S\mapsto F(x)\in S$ and $G:(x,z)\in(S\times{\cal Z})\mapsto G_{z}(x)\in S$. In addition, we assume that there exist an $x_{0}\in S$ as well as some parameters $0<\lambda(x_{0})<1$ and $c(x_{0})\geq 0$ such that
$\|F(x)-F(x_{0})\|\leq\lambda(x_{0})\,\|x-x_{0}\|+c(x_{0}),\quad\forall x\in S.$ (3.26)
To make the connection with the geometric drift condition imposed in (H1), condition (3.26) ensures that the norm is a Lyapunov function of the deterministic part of the transition.
Assume first that there exists a $\delta>0$ such that $\sup_{x\in S}\mathbb{E}\big(e^{\delta\|G_{Z}(x)\|}\big)<\infty$. Then, noting that $\|F(x)\|\leq\lambda_{0}~\|x\|+c_{\lambda}$ with $\lambda_{0}:=\lambda(x_{0})$ and $c_{\lambda}:=c(x_{0})+\|F(x_{0})\|+\lambda(x_{0})\|x_{0}\|$, we have
$\mathbb{E}\left(e^{\delta\|F(x)+G_{Z}(x)\|}\right)\leq a_{1}~e^{\delta\lambda_{0}\|x\|}\quad\mbox{\rm with}\quad a_{1}:=e^{\delta c_{\lambda}}~\sup_{x\in S}\mathbb{E}\left(e^{\delta\|G_{Z}(x)\|}\right).$
We conclude that $P(W)/W\leq\Theta\in{\cal B}_{0}(S)$ with the functions
$W(x):=\frac{1}{2}~\exp{(\delta\|x\|)}\quad\mbox{\rm and}\quad\Theta(x):=a_{1}~\exp{\left(-\delta(1-\lambda_{0})~\|x\|\right)}$
and thus, by (3.2), condition (H1) holds with $V=W\in{\cal B}_{\infty}(S)$.
Assume now that $\mathbb{E}\left(\|G_{Z}(x)\|^{p}\right)<\infty$ for some $p\geq 1$. In this scenario, letting $c_{\lambda,p}=c_{\lambda}+\sup_{x\in S}\mathbb{E}\left(\|G_{Z}(x)\|^{p}\right)^{1/p}$ we observe that
$\mathbb{E}\left(\|F(x)+G_{Z}(x)\|^{p}\right)$ $\leq\left(\mathbb{E}\left(\|G_{Z}(x)\|^{p}\right)^{1/p}+\|F(x)\|\right)^{p}\leq\left(\lambda_{0}~\|x\|+c_{\lambda,p}\right)^{p}.$
Then, by choosing $r>0$ sufficiently large so that
$\frac{c_{\lambda,p}}{\lambda_{0}r}\leq 1-\lambda_{0}\Longrightarrow\lambda_{0}\,\left(1+\frac{c_{\lambda,p}}{\lambda_{0}r}\right)\leq\lambda_{1}:=1-(1-\lambda_{0})^{2}$
we obtain that
$\mathbb{E}\left(\|F(x)+G_{Z}(x)\|^{p}\right)$ $\leq\lambda_{0}^{p}~\|x\|^{p}~\left(1+c_{\lambda,p}/(\lambda_{0}r)\right)^{p}1_{\|x\|\geq r}+\left(\lambda_{0}\,r+c_{\lambda,p}\right)^{p}$
$\leq\lambda_{1}^{p}~\|x\|^{p}+c_{\lambda,p}(r)$
where $c_{\lambda,p}(r):=\left(\lambda_{0}\,r+c_{\lambda,p}\right)^{p}$. Using the above calculations we conclude that the function $W(x):=\frac{1}{2}+\|x\|^{p}$ is such that $P(W)\leq\lambda_{1}^{p}~W+(1/2+c_{\lambda,p}(r))$, showing that (H1) holds with $V=W\in{\cal B}_{\infty}(S)$.
3.4.3 Example 1: Random iterated functions with local dissipative conditions
Consider the functions $(F,F_{z},G_{z})$ defined in Section 3.4.2 with ${\cal Z}=\mathbb{R}^{n}=S$ and
$F(x)=x+b(x)~h\quad\mbox{\rm and}\quad G_{z}(x)=z$
for some parameter $h>0$ and for a smooth function $b(x)$ with Lipschitz constant $\mbox{\rm lip}(b)>0$.
The following regularity conditions are discussed in the series of articles [6, 7, 46, 53, 54]:
$(B_{1})$ There exist $r\geq 0$, $\lambda>0$ and $c\geq 0$ such that for any $x\not=y$ we have
$\frac{(x-y)^{\top}(b(x)-b(y))}{\|x-y\|^{2}}\leq-\lambda~1_{\|x\|\vee\|y\|\geq r}+c~1_{\|x\|\vee\|y\|<r}$ (3.27)
$(B_{2})$ There exist $r\geq 0$, $\lambda>0$ and $c\geq 0$ such that for any $x\not=y$ we have
$\frac{(x-y)^{\top}(b(x)-b(y))}{\|x-y\|^{2}}\leq-\lambda~1_{\|x-y\|\geq r}+c~1_{\|x-y\|<r}$ (3.28)
The case $r=0$ is equivalent to the uniform dissipative condition. When $Z$ is a Gaussian random variable with covariance $\mathbb{E}(ZZ^{\top})=hI$ the model coincides with the transition kernel of the discretized diffusion process, under a contractivity at infinity condition (on the drift (3.27)) discussed in [53]. We point out that more general $\alpha$-stable random variables $Z$ are discussed in [46] and that condition (3.28) coincides with the local dissipative drift-type condition discussed in the articles [18, 29, 46, 54].
By applying (3.27) or (3.28) with $y=0$ we have
$\|x\|\geq r\geq r_{1}:=\frac{2\|b(0)\|}{\lambda}\Longrightarrow x^{\top}b(x)\leq-\frac{\lambda}{2}~~\|x\|^{2}$
which implies that
$x^{\top}b(x)\leq-\frac{\lambda}{2}~\|x\|^{2}+c_{r}(1)\quad\mbox{\rm with}\quad c_{r}(1)=r~\sup_{\|x\|\leq r}\|b(x)\|.$
We now let $h$ and $r$ be such that
$2~\frac{c_{1}(r)+rh\|b(0)\|}{r^{2}}\leq\lambda_{h}:=\frac{h}{2}~\left(\lambda-\mbox{\rm lip}(b)^{2}h\right)<1.$
In this case we have
$\|F(x)-F(0)\|^{2}$ $=\|x\|^{2}+2x^{\top}~(b(x)-b(0))~h+\|b(x)-b(0)\|^{2}~h^{2}$
$\leq\|x\|^{2}\left(1-(\lambda-\mbox{\rm lip}(b)^{2}h)~h+2~\frac{c_{1}(r)+rh\|b(0)\|}{r^{2}}\right)~1_{\|x\|\geq r}+c_{r}(2)$
$\leq\left(1-\lambda_{h}\right)~\|x\|^{2}+c_{r}(2)$
for some constant $c_{r}(2)>0$, implying that
$\|F(x)-F(0)\|\leq\lambda(0)~\|x\|+c(0)\quad\mbox{\rm for some}\quad 0<\lambda(0)<1\quad\mbox{\rm and}\quad c(0)>0.$
This shows that condition (3.26) is met with $x_{0}=0$ and the results of the previous subsection can be used to obtained a function $V\in{\cal B}_{\infty}(S)$ with compact level sets such that (H1) holds.
Specifically, if $Z$ is a Gaussian random variable it follows from the calculations in Section 3.4.2 that (H1) is met when $V=(1/2)\exp(\delta\|x\|)$ for some $\delta>0$. If we further assume that the function $b$ is continuous, it follows from the discussion of Section 3.4.1 that condition (H2) is met with $\kappa=\varphi_{0}$. In this context, the calculations in Section 2.6.1 show that, for any $p\geq 1$ and with $\psi_{S}(x,y)=\|x-y\|$, the following exponential contraction estimate for the Wasserstein p-distance holds:
$\mathbb{W}_{\psi_{S},p}\left(\mu_{1}P_{n},\mu_{2}P_{n}\right)\leq c\lambda^{n}\|\mu_{1}-\mu_{2}\|^{1/p}_{V},\quad\forall n\geq 0,\quad\forall\mu_{1},\mu_{2}\in\mathcal{P}_{V}(S).$
3.4.4 Example 2: A fixed point theorem
Let $F:S\rightarrow S$ be a mapping such that (3.26) holds for some $x_{0}\in S$ and parameters $\lambda(x_{0})\in]0,1[$ and $c(x_{0})\geq 0$. Next, we let $\delta>0$, $W(x)=\frac{1}{2}\exp{(\delta\|x\|)}$ and assume that there is an $r_{0}>0$ such that, for all $r\geq r_{0}$, there exists an $\alpha(r)\in]0,1[$ such that
$W(x)+W(y)\leq r\implies\|F(x)-F(y)\|\leq(1-\alpha(r))\|x-y\|.$ (3.29)
Condition (3.29) implies that $F$ is uniformly locally contractive ([16], Definition 2.1), and by a theorem of Edelstein [32] this local contraction property implies that $F$ has a unique fixed point $y^{*}\in S$ (see also Remark 2.2 and Theorem 3.5 in [16]). Below we use Corollary 2.4 to show this result, and Theorem 2.6 to prove that the sequence $(y_{n})_{n\geq 1}$ defined for some $y_{0}\in S$ by the recursion $y_{n}=F(y_{n-1})$, $n\geq 1$, converges exponentially fast to $y^{*}$.
To this aim we consider the Markov kernel $P(x,dy)=\delta_{F(x)}(dy)$. Then, following the discussion of Section 3.4.2, condition (H1) is met with $V=W$, while condition (3.29) implies that (H2) holds with the $V$-semi-distance $\kappa=\|\cdot\|$. Therefore, Corollary 2.4 ensures that $P$ has a unique invariant distribution $\mu\in\mathcal{P}_{V}(S)$. The measure $\mu$ is clearly a Dirac mass, and thus there exists a unique $y^{*}\in S$ such that
$P(y^{*},dy)=\delta_{y^{*}}(dy)\Leftrightarrow\delta_{F(y^{*})}(dy)=\delta_{y^{*}}(dy)\Leftrightarrow F(y^{*})=y^{*}$
implying that $F$ has a unique fixed point $y^{*}\in S$. Moreover, by using the first result of Theorem 2.6, it follows that there exist constants $c>0$ and $\lambda\in]0,1[$ such that
$\|y_{n}-y^{*}\|\leq c\lambda^{n}\varphi_{V}(y_{0},y^{*}),\quad\forall n\geq 0.$
3.5 Diffusion models
3.5.1 A general diffusion model
We let $S=\mathbb{R}^{n}$, $\psi_{S}(x,y)=\|x-y\|$ and we consider below the continuous time Markov semigroup $(Q_{t})_{t\in]0,\infty[}$ of a diffusion flow $y\in S\mapsto X_{t}(y)$ defined by
$dX_{t}(y)=b(X_{t}(y))\,dt+\tau(X_{t}(y))\,dB_{t}\quad\mbox{\rm with}\quad X_{0}(y)=y.$ (3.30)
In the above display, $B_{t}$ is a $n_{1}$-dimensional Brownian motion starting at the origin for some $n\geq 1$ and $b$ is a differentiable drift function from $\mathbb{R}^{n}$ into itself satisfying, for some constants $a>0$ and $c\geq 0$,
$x^{\top}b(x)\leq-a\,\|x\|^{2}+c,\quad\forall x\in S.$ (3.31)
Moreover, $\tau$ stands for some $(n\times n_{1})$-matrix valued continuous function such that, for all $x\in S$, the matrix $\tau(x)\tau(x)^{\top}$ is invertible and the map $x\mapsto\|\tau(x)\|/\|x\|$ vanishes at infinity. In what follows, for any $x\in S$ we let
$\|\tau(x)\|_{F}^{2}:=\mbox{\rm Tr}(\tau(x)\tau(x)^{\top})\leq n_{1}\|\tau(x)\|_{2}^{2}\quad\text{ with }\|\tau(x)\|_{2}^{2}=\lambda_{\max}(\tau(x)\tau(x)^{\top})$
and for all $p\geq 1$ we let $W^{p}(x):=\|x\|^{p}$.
Under the above assumptions on the functions $b$ and $\tau$, by using Itô calculus one can check that the generator $\cal G$} of the semi-group $(Q_{t})_{t\in]0,\infty[}$ is such that
${\cal G}(W^{2p})(x)$ $\leq p~W^{2p}(x)~\left(-2a+\frac{2c+\|\tau(x)\|_{F}^{2}}{\|x\|^{2}}+2(p-1)\frac{\|\tau(x)\|_{2}^{2}}{\|x\|^{2}}\right),\quad\forall p\geq 1$
with the constants $a$ and $c$ as in (3.31). Choosing $r>0$ sufficiently large so that
$\|x\|>r$ $\Longrightarrow$ $2c+\|\tau(x)\|_{F}^{2}+2(p-1)\|\tau(x)\|_{2}^{2}\leq a~\|x\|^{2}$
$\Longrightarrow$ ${\cal G}(W^{2p})(x)\leq-ap~W^{2p}(x)$~
This result shows that for all $p\geq 2$ there exists constants $a_{p}>0$ and $c_{p}\geq 0$ such that
${\cal G}(V^{p})\leq-a_{p}~V^{p}+c_{p}~~\text{ with }~~V^{p}:=\frac{1}{2}+W^{p}$ (3.32)
and thus such that (2.27) holds with $V=V^{p}$.
We now let $h\in]0,1[$, $p\geq 2$ and $P=Q_{h}$. Then, following the discussion of Section 2.6.2, the result in (3.32) implies that for $P=Q_{h}$ condition (H1) is met for $V=(1/2)+\|x\|^{p}$. We now assume that for $P=Q_{h}$ conditions (H2) is met with either $\kappa=\varphi_{0}$ or with a $V$-semi-distance $\kappa$ such that $\kappa(x,y)\geq\psi_{S}(x,y)^{p}$. In this context, it follows from (2.32) that there exist constants $c_{p}>0$ and $\lambda_{p}\in]0,1[$ such that
$\mathbb{W}_{\psi_{S},p}\left(\mu_{1}Q_{t},\mu_{2}Q_{t}\right)^{p}\leq D_{\kappa}\left(\mu_{1}Q_{t},\mu_{2}Q_{t}\right)\leq c_{p}\lambda_{p}^{t}\|\mu_{1}-\mu_{2}\|_{V},\quad\forall t\geq 0,\quad\forall\mu_{1},\mu_{2}\in\mathcal{P}_{V}(S).$
In particular, we observe that if the functions $b$ and $\tau$ are both continuous then, for $P=Q_{h}$, the local minorization condition (2.2) is met since $V\in{\cal B}_{\infty}(S)$, in which case (H2) indeed holds with $\kappa=\varphi_{0}$.
3.5.2 The overdamped Langevin diffusion model
An important class of diffusion processes arising in statistical physics and applied probability is the overdamped Langevin diffusion flow $X_{t}(y)$, defined as in (3.30) with $(n_{1},\tau(x))=(n,\sigma I)$ for some $\sigma>0$ and the drift function $b(x)=-\gamma\nabla U(x)$ associated with some parameter $\gamma>0$ and smooth potential function $U$ satisfying the following condition
$\nabla^{2}U(x)\geq\lambda~1_{\|x\|\geq r}~I-c~1_{\|x\|<r}~I\quad\mbox{for some $\lambda>0$, $c\geq 0$ and $r\geq 0$.}$ (3.33)
The local Hessian condition (3.33) is often termed “strongly convexity outside a ball”. To avoid confusions, we underline that this condition strongly differs from the strong convexity of $U$ on the non convex domain $\{x\in S:\|x\|\geq r\}$ discussed in [55, 62, 71].
In our context, we recall there exists some bounded function $\widetilde{U}$ such that $\nabla^{2}(U+\widetilde{U})\geq a~I$. For an explicit construction of such function we refer to [60]. Let $(\widetilde{Q}_{t})_{t\in]0,\infty[}$ be the continuous time Markov semigroup of the stochastic flow $\widetilde{X}_{t}(y)$ defined as $X_{t}(y)$ by replacing the drift function $b(x)=-\gamma\nabla U(x)$ by
$\widetilde{b}:=-\nabla(U+\widetilde{U})\quad\Longrightarrow\quad\nabla\widetilde{b}(x)\leq-a~I.$
Following the proof of Proposition 2.16 in [4], for any $\delta>0$ and $h>0$ there exist some constants $\delta_{h}>0$ and $c_{h}>0$ such that
$W(x):=(1/2)\exp(\delta\|x\|)\Longrightarrow\sup_{0\leq h^{\prime}\leq 1}\|\widetilde{Q}_{h^{\prime}}(W)/W\|<\infty\text{ and }\widetilde{Q}_{h}(W)/W\leq c_{h}/W^{\delta_{h}}.$ (3.34)
On the other hand, using the stochastic interpolation formula (cf. Theorem 1.2 in [23]) we check the almost sure estimate uniform estimate
$|X_{t}(x)-\widetilde{X}_{t}(x)|\leq c\quad\mbox{\rm for some finite constant $c<\infty$}.$ (3.35)
By combining (3.34) and (3.35), we obtain that
$Q_{h}(W)/W\leq e^{\delta c}~\widetilde{Q}_{h}(W)/W\leq c_{h}~e^{\delta c}~/W^{\delta_{h}}.$
Recalling that $W\in\mathcal{B}_{\infty}(S)$, we conclude that condition (3.33) ensures the existence of constants $c_{h}>0$ and $\epsilon_{h}\in]0,1[$ such that
$\begin{array}[]{l}\sup_{0\leq h^{\prime}\leq 1}\|Q_{h^{\prime}}(W)/W\|<\infty\quad\text{ and }\quad Q_{h}(W)\leq\epsilon_{h}W+c_{h}\end{array}$ (3.36)
and thus, for $P=Q_{h}$, condition (H1) is met with $V=W$. In addition, it is easily checked for $P=Q_{h}$ the local minorization condition (2.2) is met since $V\in\mathcal{B}_{\infty}(S)$, and thus for $P=Q_{h}$ condition (H2) holds with $\kappa=\varphi_{0}$. Since in addition we have $\sup_{0\leq h^{\prime}\leq 1}\|Q_{h^{\prime}}(V)/V\|<\infty$ by (3.36), the estimate (2.34) holds for any $V$-semi-distance $\phi$, as shown in Section 2.6.2. In particular, since $V$ is an exponential-type Lyapunov function, following the discussion in Section 2.6.1 there exist constants $c>0$ and $\lambda\in]0,1[$ such that
$\mathbb{W}_{\psi_{S},p}\left(\mu_{1}Q_{t},\mu_{2}Q_{t}\right)^{p}\leq c~\lambda^{t}~\|\mu_{1}-\mu_{2}\|_{V},\quad\forall t\geq 0,\quad\forall p\geq 1,\quad\forall\mu_{1},\mu_{2}\in\mathcal{P}_{V}(S)$
where for all $p\geq 1$ the constant $c_{p}$ is as defined in (2.25).
4 Proofs
4.1 Proof of Lemma 1.1
Proof.
We clearly have the estimate
$\beta_{\psi,\phi}(P)\geq\ell_{\psi,\phi}^{\star}(P):=\sup_{(x,y)\in S^{2}~:~\psi(x,y)>0}\ell_{\psi,\phi}(P)(x,y)$
and thus, recalling that since $\psi$ is a semi-distance we have $\psi(x,y)>0$ if and only if $x\neq y$, to prove the lemma it remains to show that
$\beta_{\psi,\phi}(P)\leq\ell_{\psi,\phi}^{\star}(P).$ (4.1)
To this aim, following [72], page 67, we remark that since $\phi$ is l.s.c. we have $\phi(x,y)=\lim_{k\rightarrow\infty}\phi_{k}(x,y)$ for all $x,y\in S$, where for all $k\geq 1$ the function $\phi_{k}$ is defined by
$\phi_{k}(x,y)=\inf_{x^{\prime},y^{\prime}\in S}\Big\{\min\Big(\phi(x^{\prime},y^{\prime}),k\Big)+k\Big(\psi_{S}(x,x^{\prime})+\psi_{S}(y,y^{\prime})\Big)\Big\},\quad x,y\in S.$ (4.2)
It is easy to verify that $(\phi_{k})_{k\geq 1}$ is a non-decreasing sequence of bounded and continuous functions such that $0\leq\phi_{k}(x,y)\leq\min(\phi(x,y),k)$ for all $k\geq 1$ and all $x,y\in S$.
As a first step towards proving (4.1) we show that
$D_{\phi}(\mu_{1},\mu_{2})=\limsup_{k\rightarrow\infty}D_{\phi_{k}}(\mu_{1},\mu_{2}),\quad\forall\mu_{1},\mu_{2}\in\mathcal{P}_{V}(S).$ (4.3)
To do so we let $\mu_{1},\mu_{2}\in\mathcal{P}_{V}(S)$ be arbitrary and remark that the set $\Pi(\mu_{1},\mu_{2})$ is compact w.r.t. the topology of the weak convergence; this follows from Lemma 4.4 in [72] and Prokhorov’s theorem. Next, for any $k\geq 1$ we let $\pi_{k}\in\Pi(\mu_{1},\mu_{2})$ be such that $D_{\phi_{k}}(\mu_{1},\mu_{2})=\pi_{k}(\phi_{k})$; remark that such a measure $\pi_{k}$ exists since $\phi_{k}$ is a $V$-semi-distance and $S$ is Polish. Since the set $\Pi(\mu_{1},\mu_{2})$ is compact, the sequence $(\pi_{k})_{k\geq 1}$ admits a subsequence $(\pi_{k_{i}})_{i\geq 1}$ that converges weakly to some measure $\tilde{\pi}\in\Pi(\mu_{1},\mu_{2})$. Then, noting that $\phi_{l}\leq\phi_{k}$ for all $k\geq l\geq 1$, and using the fact that the functions $(\phi_{l})_{l\geq 1}$ are continuous and bounded, for any fixed $l\geq 1$ we have
$\tilde{\pi}(\phi_{l})=\lim_{i\rightarrow\infty}\pi_{k_{i}}(\phi_{l})\leq\limsup_{i\rightarrow\infty}\pi_{k_{i}}(\phi_{k_{i}})=\limsup_{i\rightarrow\infty}D_{\phi_{k_{i}}}(\mu_{1},\mu_{2})\leq\limsup_{k\rightarrow\infty}D_{\phi_{k}}(\mu_{1},\mu_{2}).$ (4.4)
On the other hand, we have
$\limsup_{k\rightarrow\infty}D_{\phi_{k}}(\mu_{1},\mu_{2})\leq D_{\phi}(\mu_{1},\mu_{2})\leq\tilde{\pi}(\phi)=\lim_{l\rightarrow\infty}\tilde{\pi}(\phi_{l})$ (4.5)
where the first inequality holds since $\phi_{k}\leq\phi$ for all $k\geq 1$, the second inequality holds since $\tilde{\pi}\in\Pi(\mu_{1},\mu_{2})$, while the equality holds by the monotone convergence theorem. Then, combining (4.4) and (4.5) shows (4.3).
We now let $k\geq 1$ be arbitrary. Then, since the $V$-semi-distance $\phi_{k}$ is continuous it follows from Corollary 5.22 in [72] that there exists a Markov transition $Q_{k}$ on $S^{2}$ such that, for all $x,y\in S$, the measure $\pi^{(k)}_{x,y}:=Q_{k}((x,y),d(u,v))$ belongs to $\Pi(\delta_{x}P,\delta_{y}P)$ and satisfies $\pi^{(k)}_{x,y}(\phi_{k})=D_{\phi_{k}}(\delta_{x}P,\delta_{y}P)$.
We now show that under the assumption of the lemma we have
$D_{\phi_{k}}(\delta_{x}P,\delta_{y}P)=\int~\phi_{k}(u,v)Q_{k}((x,y),d(u,v))\leq~\ell_{\psi,\phi}^{\star}(P)\psi(x,y),\quad\forall x,y\in S.$ (4.6)
Since $\phi_{k}\leq\phi$, the inequality in (4.6) clearly holds for any $x,y\in S$ such that $\psi(x,y)>0$. Let $x,y\in S$ be such that $\psi(x,y)=0$. Then, since $\psi$ is a semi-distance, we have $x=y$. On the other hand, since $\phi$ is a also a semi-distance, it follows from (4.2) that $\phi_{k}$ is a semi-distance, and thus we have the implication $x=y\implies D_{\phi_{k}}(\delta_{x}P,\delta_{y}P)=0$. This shows that $\psi(x,y)=0\implies D_{\phi_{k}}(\delta_{x}P,\delta_{y}P)=0$, and thus the inequality in (4.6) also holds when $\psi(x,y)=0$. This concludes the proof of (4.6).
To proceed further we remark that since $\psi$ is l.s.c. there exists a measure $\pi\in\Pi(\mu_{1},\mu_{2})$ such that $D_{\psi}(\mu_{1},\mu_{2})=\pi(\psi)$, and we also remark that
$(\pi Q_{k})(d(u,v)):=\int~\pi(d(x,y))Q_{k}((x,y),d(u,v))\in\Pi(\mu_{1}P,\mu_{2}P).$ (4.7)
Consequently, by using first (4.7) and then by integrating both side of (4.6) using the measure $\pi$ we have just defined, we obtain that
$D_{\phi_{k}}(\mu_{1}P,\mu_{2}P)$ $\leq(\pi Q_{k})(\phi_{k})\leq\ell_{\psi,\phi}^{\star}(P)\pi(\psi)=\ell_{\psi,\phi}^{\star}(P)D_{\psi}(\mu_{1},\mu_{2})$ (4.8)
showing that it $D_{\psi}(\mu_{1},\mu_{2})>0$ then
$\frac{D_{\phi_{k}}(\mu_{1}P,\mu_{2}P)}{D_{\psi}(\mu_{1},\mu_{2})}\leq\ell_{\psi,\phi}^{\star}(P).$
Together with (4.3), this shows (4.1) and the proof of the lemma is complete.
Remark 4.1.
Under the additional assumption that $\phi$ is continuous, by Corollary 5.22 in [72] there exists a Markov transition $Q$ on $S^{2}$ such that
$D_{\phi}(\delta_{x}P,\delta_{y}P)=~\int~\phi(u,v)~Q((x,y),d(u,v)),\quad\forall x,y\in S.$ (4.9)
In this case, (4.1) directly follows by applying (4.8) with $Q_{k}$ replaced by $Q$. Assuming only that $\phi$ is l.s.c., (4.3) implies that the mapping $(x,y)\mapsto D_{\phi}(\delta_{x}P,\delta_{y}P)$ is measurable but it is unclear whether or not there exists a Markov kernel $Q$ such that (4.9) holds.
4.2 Proof of (1.13)
Note that
$\begin{array}[]{l}\sup{\{\mbox{\rm osc}_{V}(K(f))~:~f\in{\cal B}(S)~;~\mbox{\rm osc}_{W}(f)\leq 1\}}\\ \\ \displaystyle=\sup_{(x,y)\in S^{2}}\frac{1}{\varpi_{V}(x,y)}~\sup{\{|(\delta_{x}K)(f)-(\delta_{y}K)(f)|~:~f\in{\cal B}(S)~;~\mbox{\rm osc}_{W}(f)\leq 1\}}\\ \\ \displaystyle=\sup_{(x,y)\in S^{2}}\frac{\|\delta_{x}K-\delta_{y}K\|_{W}}{\varpi_{V}(x,y)}=\sup_{(x,y)\in S^{2}~:~x\neq y}\ell_{\varphi_{V},\varphi_{W}}(K)(x,y).\end{array}$
On the other hand, by (1.8) we check that
$\begin{array}[]{l}\|\mu_{1}K-\mu_{2}K\|_{W}\\ \\ =\sup{\left\{\mbox{\rm osc}_{V}(K(f))\times|(\mu_{1}-\mu_{2})[K(f)/\mbox{\rm osc}_{V}(K(f))]|~:~f\in{\cal B}(S)~;~\mbox{\rm osc}_{W}(f)\leq 1\right\}}\\ \\ \leq\sup{\{\mbox{\rm osc}_{V}(K(f))~:~f\in{\cal B}(S)~;~\mbox{\rm osc}_{W}(f)\leq 1\}}\times\|\mu_{1}-\mu_{2}\|_{V}\end{array}$
This implies that
$\beta_{\varphi_{V},\varphi_{W}}(K)\leq\sup_{(x,y)\in S^{2}~:~x\neq y}\ell_{\varphi_{V},\varphi_{W}}(K)(x,y)$
The reverse inequality is obtained choosing Dirac measures $(\mu_{1},\mu_{2})=(\delta_{x},\delta_{y})$ in (1.11). This ends the proof of (1.13).
4.3 Proof of Lemma 2.2
The first statement in (2.9) is trivial. To prove the second statement in (2.9) let $x,y\in S$ be such that $\varphi(x,y)>0$ and $\pi_{x,y}\in\Pi(\delta_{x}P,\delta_{y}P)$ be such that $D_{\varphi}(\delta_{x}P,\delta_{y}P)=\pi_{x,y}(\varphi)$. Then, by Lemma 1.1, we have
$D_{\varphi}(\delta_{x}P,\delta_{y}P)=\pi_{x,y}(\varphi)\leq\beta_{\varphi}(P)~\varphi(x,y).$
By Jensen’s inequality, for any $\iota\in]0,1[$ we have
$D_{\varphi^{\iota}}(\delta_{x}P,\delta_{y}P)\leq\pi_{x,y}(\varphi^{\iota})\leq\pi_{x,y}(\varphi)^{\iota}=D_{\varphi}(\delta_{x}P,\delta_{y}P)^{\iota}\leq\beta_{\varphi}(P)^{\iota}~\varphi(x,y)^{\iota}$
implying that $D_{\varphi^{\iota}}(\delta_{x}P,\delta_{y}P)\leq\beta_{\varphi}(P)^{\iota}~\varphi(x,y)^{\iota}$ which, by Lemma 1.1, implies that we have $\beta_{\varphi^{\iota}}(P)\leq\beta_{\varphi}(P)^{\iota}$. This ends the proof of (2.9).
To show (2.10) let $\psi$ be a $V$-semi-distance such that $a\,\varphi\leq\psi\leq b\,\varphi$ for some constants $0<a\leq b$. When $\beta_{\psi}(P)=\infty$ the estimate (2.10) is immediate. When $\beta_{\psi}(P)<\infty$ for any $\mu_{1},\mu_{2}\in\mathcal{P}_{V}(S)$ we have
$a\,D_{\varphi}(\mu_{1}P,\mu_{2}P)\leq D_{\psi}(\mu_{1}P,\mu_{2}P)\leq\beta_{\psi}(P)~D_{\psi}(\mu_{1},\mu_{2})\leq b\,\beta_{\psi}(P)~D_{\varphi}(\mu_{1},\mu_{2})$
which implies that $D_{\varphi}(\mu_{1}P,\mu_{2}P)\leq(b/a)\,\beta_{\psi}(P)~D_{\varphi}(\mu_{1},\mu_{2})$. The result in (2.10) follows.
To show (2.11) we let $\psi$ be a $V$-semi-distance such that $\varphi\leq\psi$ and such that $\beta_{\psi}(P)<1$. Next, we let $x,y\in S$ be such that $\psi(x,y)>0$ and $\pi_{x,y}\in\Pi(\delta_{x}P,\delta_{y}P)$ be such that $D_{\psi}(\delta_{x}P,\delta_{y}P)=\pi_{x,y}(\psi)$. Then,
$D_{\varphi}(\delta_{x}P,\delta_{y}P)\leq\pi_{x,y}(\varphi)\leq\pi_{x,y}(\psi)=D_{\psi}(\delta_{x}P,\delta_{y}P)\leq\beta_{\psi}(P)\,\psi(x,y)$ (4.10)
where the last inequality holds by Lemma 1.1. Using (4.10), it follows that for any $\iota>0$ we have
$\frac{D_{\iota\varphi+\psi}(\delta_{x}P,\delta_{y}P)}{\iota\varphi(x,y)+\psi(x,y)}\leq\frac{\pi_{x,y}(\iota\varphi+\psi)}{\iota\varphi(x,y)+\psi(x,y)}\leq(1+\iota)~\beta_{\psi}(P)~\frac{\psi(x,y)}{\iota\varphi(x,y)+\psi(x,y)}$
and thus, for any $\delta\in]0,1[$, by Lemma 1.1 we have the implication:
$\beta_{\psi}(P)<\frac{\delta}{1+\iota}\Longrightarrow\beta_{\iota\varphi+\psi}(P)\leq(1+\iota)~\beta_{\psi}(P)\leq\delta$ (4.11)
showing (2.11). Lastly, noting that for $\iota\leq 1-\beta_{\psi}(P)$ we have $(1+\iota)\beta_{\psi}(P)\leq 1-(1-\beta_{\psi}(P))^{2}$, it follows from (4.11) that
$\beta_{\iota\varphi+\psi}(P)\leq 1-(1-\beta_{\psi}(P))^{2}$
showing (2.12). This ends the proof of the lemma.
4.4 Proof of Theorem 2.3
The contraction estimate for the weighted discrete metric $\varphi_{\rho}$ stated in (2.14) is known, but for sake of completeness a detailed proof is provided below in Section 4.4.1. The estimate (2.15) is a consequence of Lemma 2.2 combined with (2.14). To see this let $r>r_{\epsilon}\vee r_{0}$ so that $\beta:=\beta_{\varphi_{\rho_{\epsilon}(r)}}(P)\in]0,1[$ by (2.14), and let $\phi$ be a $V$-semi-distance such that $\|\phi/\varphi_{V}\|\leq(1-\beta)\rho_{\epsilon}(r)$. Then, using (2.5) for the third inequality,
$\phi\leq\|\phi/\varphi_{V}\|\varphi_{V}\leq(1-\beta)\rho_{\epsilon}(r)\,\varphi_{V}\leq(1-\beta)\varphi_{\rho_{\epsilon}(r)}$
and (2.15) follows by applying (2.12) with $\iota=\|\phi/\varphi_{\rho_{\epsilon}(r)}\|$, $\varphi=\phi/\iota$ and $\psi=\varphi_{\rho_{\epsilon}(r)}$. The proof of (2.13) is presented below in Section 4.4.2.
4.4.1 Proof of (2.14)
Let $x,y\in S$ be such that $x\neq y$ and $r>r_{\epsilon}\vee r_{0}$ be fixed in what follows. Then, under the assumption that (H2) holds with the semi-distance $\kappa=\varphi_{0}$, we have
$\|\delta_{x}P-\delta_{y}P\|_{\tiny tv}\leq 1-\alpha(r)<1.$ (4.12)
In what follows the functions $(\varpi_{\rho},\varphi_{\rho})$ indexed by $\rho>0$ are as defined in (2.3) and (2.4).
Assume first that $\varpi_{V}(x,y)\geq r$. In this scenario, for all $\rho>0$ we have
$\begin{split}\ell_{\varphi_{\rho}}(P)(x,y)&\leq\ell_{\varpi_{\rho}}(P)(x,y)=\frac{1+\rho P(V)(x)+\rho P(V)(y)}{1+\rho\varpi_{V}(x,y)}\\ &\leq\frac{1+\rho\big(1+\epsilon\varpi_{V}(x,y)\big)}{1+\rho\varpi_{V}(x,y)}\\ &=\epsilon+\frac{1+\rho-\epsilon}{1+\rho r}=1-(1-\epsilon)~\frac{\rho r}{1+\rho r}~\left(1-\frac{r_{\epsilon}}{r}\right)\end{split}$ (4.13)
where the first inequality uses the assumption that (H1) holds with $c=1/2$ and the last inequality uses (4.12). Simple calculations show that
$(1-\epsilon)\frac{r\rho_{\epsilon}(r)}{1+r\rho_{\epsilon}(r)}\left(1-\frac{r_{\epsilon}}{r}\right)$ $=~\frac{\alpha(r)}{2}~\frac{1-\epsilon}{(1+\epsilon)+\alpha(r)/2}~\left(1-\frac{r_{\epsilon}}{r}\right)\geq~\frac{\alpha(r)}{2}~\delta_{\epsilon}(r)$
and thus we have the implication
$\varpi_{V}(x,y)\geq r\implies\ell_{\varphi_{\rho_{\epsilon}(r)}}(P)(x,y)\leq\ell_{\varpi_{\rho_{\epsilon}(r)}}(P)(x,y)\leq 1-\delta_{\epsilon}(r)\frac{\alpha(r)}{2}.$ (4.14)
In the reverse angle, if $\varpi_{V}(x,y)\leq r$ then, following similar calculations as in (4.13),
$\ell_{\varphi_{\rho}}(P)(x,y)$ $\leq\frac{1-\alpha(r)+\rho\big(1+\epsilon\varpi_{V}(x,y)\big)}{1+\rho\varpi_{V}(x,y)}\leq 1-\alpha(r)+\rho+r\epsilon\rho,\quad\forall\rho>0$
from which we readily obtain that
$\varpi_{V}(x,y)\leq r\implies\ell_{\varphi_{\rho_{\epsilon}(r)}}(P)(x,y)\leq 1-\frac{\alpha(r)}{2}\leq 1-\delta_{\epsilon}(r)\frac{\alpha(r)}{2}.$ (4.15)
Combining (4.14) and (4.15) show that
$\ell_{\varphi_{\rho_{\epsilon}(r)}}(P)(x,y)\leq 1-\delta_{\epsilon}(r)\frac{\alpha(r)}{2}$
and the result in (2.14) follows from Lemma 1.1.
4.4.2 Proof of (2.13)
In what follows the functions $(\varpi_{\rho},\varphi_{\rho})$ indexed by $\rho>0$ are as defined in (2.3) and (2.4), and we extend the definition of the function $\ell_{\phi}(P):(x,y)\in S^{2}\rightarrow[0,\infty]$ defined in Section 1.3 for a $V$-semi-distance $\phi$ to any l.s.c. function $\phi:S^{2}\rightarrow[0,\infty)$.
We start with some preliminary calculations. We let $\rho>0$, $\iota\in]0,1[$ and $x,y\in S$ be such that $x\neq y$. Next, we let $\pi_{x,y}\in\Pi(\delta_{x}P,\delta_{y}P)$ be such that $\pi_{x,y}(\kappa)=D_{\kappa}(\delta_{x}P,\delta_{y}P)$, so that $\ell_{\kappa}(P)(x,y)=\pi_{x,y}(\kappa)/\kappa(x,y)$. In addition, noting that
$\pi_{x,y}\left(\varpi_{\rho}\right)=1+\rho\,P(V)(x)+\rho\,P(V)(y)=D_{\varpi_{\rho}}(\delta_{x}P,\delta_{y}P)$ (4.16)
it follows that
$\ell_{\varpi_{\rho}}(P)(x,y)=\frac{\pi_{x,y}\left(\varpi_{\rho}\right)}{\varpi_{\rho}(x,y)}.$
Using the above expressions for $\ell_{\kappa}(P)(x,y)$ and for $\ell_{\varpi_{\rho}}(P)(x,y)$, as well as Hölder inequality, we obtain that
$\begin{split}\ell_{\kappa_{\iota,\rho}}(P)(x,y)\leq\frac{\pi_{x,y}(\kappa_{\iota,\rho})}{\varpi_{\rho}(x,y)^{1-\iota}\kappa(x,y)^{\iota}}&\leq\ell_{\varpi_{\rho}}(P)(x,y)^{1-\iota}~\ell_{\kappa}(P)(x,y)^{\iota}\\ &\leq\ell_{\varpi_{\rho}}(P)(x,y)^{1-\iota}\end{split}$ (4.17)
where the last inequality uses the fact that $\beta_{\kappa}(P)\leq 1$ under (H2).
To prove (2.13) let $x,y\in S$ be such that $x\neq y$ and $r>r_{\epsilon}\vee r_{0}$ be fixed in what follows. Then,
$\begin{split}\varpi_{V}(x,y)\geq r\Longrightarrow\ell_{\kappa_{\iota,\rho_{\epsilon}(r)}}(P)(x,y)\leq\ell_{\varpi_{\rho_{\epsilon}(r)}}(P)(x,y)^{1-\iota}\leq\displaystyle\Big(1-\delta_{\epsilon}(r)\frac{\alpha(r)}{2}\Big)^{1-\iota}\end{split}$ (4.18)
where the first inequality holds (4.17) while the second inequality holds by (4.14).
In the reverse angle, if $\varpi_{V}(x,y)\leq r$ we have, under the assumption that (H1) holds with $c=1/2$,
$\ell_{\varpi_{\rho_{\epsilon}(r)}}(P)(x,y)\leq 1+\rho_{\epsilon}(r)(1+\epsilon\,r)\leq 1+\alpha(r)$
where the second inequality holds since $\rho_{\epsilon}(r)\leq\alpha(r)/(1+\epsilon r)$. Consequently,
$\begin{split}\varpi_{V}(x,y)\leq r\Longrightarrow\ell_{\kappa_{\iota,\rho_{\epsilon}(r)}}(P)(x,y)&\leq\ell_{\kappa}(P)(x,y)^{\iota}~\ell_{\varpi_{\rho_{\epsilon}(r)}}(P)(x,y)^{1-\iota}\\ &\leq(1-\alpha(r))^{\iota}~(1+\alpha(r))^{1-\iota}\\ &\leq(1-\alpha(r)^{2})^{1-\iota}\end{split}$ (4.19)
where first inequality holds (4.17) and where the last inequality comes from the fact that, for any $1/2\leq\iota<1$, we have
$(1-\alpha(r))^{\iota}~(1+\alpha(r))^{1-\iota}=(1-\alpha(r))^{\iota-(1-\iota)}~(1-\alpha(r)^{2})^{1-\iota}\leq(1-\alpha(r)^{2})^{1-\iota}.$
Combining (4.18) and (4.19) show that
$\ell_{\kappa_{\iota,\rho_{\epsilon}(r)}}(P)(x,y)\leq\left(1-\alpha(r)\left(\alpha(r)\wedge\frac{\delta_{\epsilon}(r)}{2}\right)\right)^{1-\iota}$
and the result in (2.13) follows from Lemma 1.1.
4.5 Proof of Theorem 2.6
We start by proving the first part of the theorem. Since for any $V$-semi-distance $\kappa$ we have $\beta_{\varphi_{V},\kappa}(P_{n})\leq\|\kappa/\varphi_{V}\|~\beta_{\kappa}(P_{n})$ for all $n\geq 1$ the result when $\iota<1$ follows from the first part of Theorem 2.5. Hence, to prove the first part of the theorem it remains to consider the case $\iota=1$.
To this aim, let $\kappa$ be a $V$-semi-distance such that (H2) holds and remark that, using (2.8) and the remark made just before the statement of Theorem 2.3, we can without loss of generality assume in what follows that (H1) holds with $c=1/2$. Next, let $c_{\kappa}=\|\kappa/\varpi_{V}\|<\infty$ and $r_{1}>r_{\epsilon}\vee r_{0}$, and note that for all $x,y\in S$ we have
${\kappa(x,y)}/{c_{\kappa}}\leq\varpi_{V}(x,y)$ (4.20)
To proceed further let $\tilde{\kappa}:=(1/c_{\kappa})\kappa$. We remark that $\tilde{\kappa}$ is a $V$-semi-distance that also satisfies (H2), and we let $\kappa^{\prime}:=(\varpi_{V}\tilde{\kappa})^{1/2}$. Then, by using the first result stated in Theorem 2.5, it follows that there exist constants $c_{0}>0$ and $\lambda\in]0,1[$ such that $\beta_{\kappa^{\prime}}(P_{n})\leq c_{0}\lambda^{n}$ for all $n\geq 0$. On the other hand, using (4.20) we have $\tilde{\kappa}\leq\kappa^{\prime}\leq\varphi_{V}$, where the last inequality uses the fact that $\kappa(x,x)=0$ for all $x\in S$. Consequently, for all $n\geq 0$ and $\mu_{1},\mu_{2}\in\mathcal{P}_{V}(S)$ we have
$D_{\tilde{\kappa}}(\mu_{1}P_{n},\mu_{2}P_{n})\leq D_{\kappa^{\prime}}(\mu_{1}P_{n},\mu_{2}P_{n})$ $\leq\beta_{\kappa^{\prime}}(P_{n})D_{\kappa^{\prime}}(\mu_{1},\mu_{2})\leq c_{0}~\lambda^{n}~D_{\varphi_{V}}(\mu_{1},\mu_{2})$
and thus, recalling that $\tilde{\kappa}=(1/c_{\kappa})\kappa$, we obtain that
$D_{\kappa}\big(\mu_{1}P_{n},\mu_{2}P_{n}\big)\leq c_{0}\,c_{\kappa}\lambda^{n}~D_{\varphi_{V}}(\mu_{1},\mu_{2}),\quad\forall n\geq 0,\quad\forall\mu_{1},\mu_{2}\in\mathcal{P}_{V}(S)$ (4.21)
showing the first part of the theorem follows.
To show the second part of the theorem we assume that $\beta_{\kappa,\varphi_{0}}(P)<\infty$. In this context, there exists a constant $c>0$ such that
$D_{\varphi_{0}}(\delta_{x}P_{n},\delta_{y}P_{n})\leq cD_{\kappa}(\delta_{x}P_{n-1},\delta_{y}P_{n-1}),\quad\forall n\geq 2,\quad\forall x,y\in S.$ (4.22)
We now let $r_{1}>r_{\epsilon}$ be arbitrary. Then, by using (4.21) and (4.22), there exist constants $\alpha\in]0,1]$ and $m\in\mathbb{N}$ such that the following implication holds:
$\varphi_{V}(x,y)\leq r_{1}\implies D_{\varphi_{0}}(\delta_{x}P_{r},\delta_{y}P_{r})\leq(1-\alpha)\varphi_{0}(x,y).$ (4.23)
By using (2.8) and by replacing $V$ by $\overline{V}:=(1-\epsilon)/2+\big(\epsilon(1-\epsilon)/(2c)\big)V$ if necessary, we can without loss of generality assume that (H1) holds with $c=(1-\epsilon)/2$, in which case
$P_{m}(V)\leq\epsilon V+\frac{1}{2}.$ (4.24)
Then, using (4.23) and (4.24), we can repeat word-for-word the calculations in the proof of (2.14), replacing $P$ by $P_{m}$, $r$ by $r_{1}$ and $\alpha(r)$ by $\alpha$, to obtain that there exist constants $\rho>0$ and $\lambda\in]0,1[$ such that $\beta_{\varphi_{\rho}}(P_{m})<\lambda$. Then, repeating word-by-word the proof of (2.20), it follows that there exist constants $c>0$ and $\lambda\in]0,1[$ such that $\beta_{\varphi_{V}}(P_{nm})\leq c\lambda^{n}$ for all $n\geq 0$, and since $\beta_{\varphi_{V}}(P)<\infty$ under (H1), we obtain that $\beta_{\varphi_{V}}(P_{nm})\leq c^{\prime}\lambda^{n}$ for all $n\geq 0$ and some constant $c^{\prime}>0$. Consequently, for any $V$-semi-distance function $\phi$ and measures $\mu_{1},\mu_{2}\in\mathcal{P}_{V}(S)$, we have
$D_{\phi}\big(\mu_{1}P_{n},\mu_{2}P_{n}\big)\leq\|\phi/\varphi_{V}\|D_{\varphi_{V}}\big(\mu_{1}P_{n},\mu_{2}P_{n}\big)\leq\|\phi/\varphi_{V}\|c^{\prime}\lambda^{n}D_{\varphi_{V}}(\mu_{1},\mu_{2}),\quad\forall n\geq 0.$
This ends the proof of the theorem.
4.6 Proof of (2.24) and (2.25)
For functions of the form $V(x)=1/2+\|x\|^{p}$ for some $p>0$ we have
$\frac{1}{2^{(p-1)_{+}}}~\|x-y\|^{p}\leq\|x\|^{p}+\|y\|^{p}\leq\varpi_{V}(x,y)$ (4.25)
This ends the proof of (2.24). For functions of the form $V(x)=1+e^{\delta\|x\|}$, for any $q\geq 1$ we have
$\varpi_{e}(x,y)\leq 2\left(e^{\delta(\|x\|+\|y\|)/2}-1\right)\leq 2~e^{\delta\|x\|/2}~e^{\delta\|y\|/2}\leq\varpi_{V}(x,y)$ (4.26)
This ends the proof of (2.25).
4.7 Proof of (3.7) and (3.9)
Note that
$V(x)\geq V(1/2)=2^{1-\iota}>0\Longrightarrow V\in{\cal B}_{0}(S)$
For any $\iota\in]0,1/2[$ there exists some finite constant $c>0$ such that
$P(V)(x)\leq\epsilon_{\iota}~V(x)+c\quad\mbox{\rm with}\quad\epsilon_{\iota}:=\frac{1}{2(1-\iota)}<1$
We check this claim using the formula
$2(1-\iota)~P(V)(x)$ $=$ $\frac{1}{x}\left(x^{1-\iota}+\left(1-(1-x)^{1-\iota}\right)\right)+\frac{1}{1-x}\left((1-x)^{1-\iota}+\left(1-x^{1-\iota}\right)\right)$
$=$ $V(x)+f(x)+f(1-x)$
with the uniformly bounded function
$\begin{array}[]{l}\displaystyle f(x):=\frac{1}{x}\left(1-(1-x)^{1-\iota}\right)=\frac{1}{(1-x)^{\iota}}+\frac{1}{x}~\left(1-\frac{1}{(1-x)^{\iota}}\right)\leq\frac{1}{(1-x)^{\iota}}\wedge\frac{1}{x}\\ \\ \displaystyle\Longrightarrow\forall\delta\in]0,1[\qquad 0\leq f(x)\leq\frac{1}{\delta^{\iota}}~1_{\delta\leq 1-x<1}+\frac{1}{1-\delta}~1_{1-\delta<x<1}\end{array}$
This ends the proof of (3.7).
To prove (3.9), consider the parameters $\iota\in]0,1[$ such that
$\iota\leq 1\wedge\gamma\quad\mbox{\rm and}\quad 0<\frac{\iota}{2}<\frac{1-\delta}{1+(1-\delta)}~\left(\leq\frac{1}{2}\right)$
Observe that
$\frac{1}{x}~\int_{0}^{x}~\left(\frac{1}{y^{\iota}}+y\right)~dy=\frac{1}{1-\iota}~\frac{1}{x^{\iota}}+\frac{x}{2}$
We also check that
$\int_{x}^{\infty}\left(\frac{1}{y^{\iota}}+y\right)\sqrt{\frac{2}{\pi}}~e^{-(y-x)^{2}/2}dy\leq\frac{1}{x^{\iota}}+x+c_{0}\quad\mbox{\rm with}\quad c_{0}:=\int_{0}^{\infty}z\sqrt{\frac{2}{\pi}}~e^{-z^{2}/2}dz$
$P(V)(x)$ $\leq$ $\frac{\delta}{2}\left(\frac{1}{1-\iota}~\frac{1}{x^{\iota}}+\frac{x}{2}+\frac{1}{x^{\iota}}+x+c_{0}\right)+(1-\delta)~\nu(V)$
$\leq$ $\frac{\delta}{2}~\left(1+\frac{1}{1-\iota}\right)~\frac{1}{x^{\iota}}+\frac{\delta}{2}~\left(1+\frac{1}{2}\right)~x+c_{1}\quad\mbox{\rm with}\quad c:=\frac{\delta}{2}~c_{0}+(1-\delta)~\nu(V)$
$\leq$ $\epsilon~V(x)+c\quad\mbox{\rm with}\quad\epsilon:=\frac{\delta}{2}~\left(1+\frac{1}{1-\iota}\right)<1$
This ends the proof of (3.9).

### v2

@@
 Specifically, the quantity $D_{\psi_{S}}(\mu_{1},\mu_{2})$ is the Kantorovitch distance between $\mu_{1}$ and $\mu_{2}$ (a.k.a.
-the Wasserstein 1-distance) while for $p\geq 1$ the Wasserstein $p$-distance $\mathbb{W}_{\psi_{S},p}(\mu_{1},\mu_{2})$ is defined by $\mathbb{W}_{\psi_{S},p}(\mu_{1},\mu_{2})=D_{\psi_{S}^{p}}(\mu_{1},\mu_{2})^{p}$.
+the Wasserstein 1-distance) while for $p\geq 1$ the Wasserstein $p$-distance $\mathbb{W}_{\psi_{S},p}(\mu_{1},\mu_{2})$ is defined by $\mathbb{W}_{\psi_{S},p}(\mu_{1},\mu_{2})^{p}=D_{\psi_{S}^{p}}(\mu_{1},\mu_{2})$.
 We also have the equality $D_{\varphi_{0}}(\mu_{1},\mu_{2})=\|\mu_{1}-\mu_{2}\|_{tv}$, that is $D_{\varphi_{0}}(\mu_{1},\mu_{2})$ is the total variation distance between the probability measures $\mu_{1}$ and $\mu_{2}$, and the weighted total variation norm $\|\mu_{1}-\mu_{2}\|_{V}$ is related to the Kantorovich semi-distance (1.6) by the formula
@@
 Under this assumption on $\phi$, for any $\mu_{1},\mu_{2}\in{\cal P}_{V}(S)$ we indeed have
-$D_{\phi}(\mu_{1},\mu_{2})\leq(\mu_{1}\otimes\mu_{2})(\phi)\leq\mu_{1}(V)+\mu_{2}(V)<\infty$
+$D_{\phi}(\mu_{1},\mu_{2})\leq(\mu_{1}\otimes\mu_{2})(\phi)\leq\|\phi/\varpi_{V}\|\left(\mu_{1}(V)+\mu_{2}(V)\right)<\infty$
 with the tensor product measure $(\mu_{1}\otimes\mu_{2})(d(x,y)):=\mu_{1}(dx)\mu_{2}(dy)$.
@@
 In many applications $P=KL$ for some Markov integral operators $K$ and $L$ (see Section 3.3). In this context, if there exists a l.s.c. function $W$, lower bounded away from zero and such that
-$K(W)\leq\epsilon~V+c\quad\mbox{\rm and}\quad L(V)\leq\epsilon~W+c$ (2.1)
-for some constants $\epsilon_{0}\in]0,1[$ and $c_{0}\geq 0$, then $P$ is a Markov $V$-operator and (H1) holds with $\epsilon=\epsilon_{0}^{2}$ and $c=(1+\epsilon_{0})c_{0}$.
+$K(W)\leq\epsilon_{1}~V+c_{1}\quad\mbox{\rm and}\quad L(V)\leq\epsilon_{2}~W+c_{2}$ (2.1)
+for some constants $c_{1},c_{2}\geq 0$ and $\epsilon_{1},\epsilon_{2}\in]0,1[$, then $P$ is a Markov $V$-operator and (H1) holds with $\epsilon=\epsilon_{1}\epsilon_{2}$ and $c=(c_{1}\epsilon_{2}+c_{2})$.
 In addition to (H1), the results presented in this section will assume that the following additional condition on $P$ is satisfied:
@@
 with the semi-distance $\kappa_{\iota,V}$ as defined in (1.5).
-Moreover, if $\beta_{\kappa,\varphi_{0}}(P)<\infty$ then, for any $V$-semi-distance $\phi$, there exists a constant $c^{\prime}\geq 0$ such that
+If, in addition, we have $\beta_{\kappa,\varphi_{0}}(P)<\infty$, then, for any $V$-semi-distance $\phi$, there exists a constant $c^{\prime}\geq 0$ such that
 $\beta_{\varphi_{V},\phi}(P_{n})\leq c^{\prime}~\lambda^{n},\quad\forall n\geq 0.$
@@
 […]
-It is also worth noting that the second part of Theorem 2.6 implies that if (H2) holds for the discrete metric (that is with $\kappa=\varphi_{0}$) then the $V$-Dobrushin contraction coefficient $\beta_{\varphi_{V}}(P_{n})$ converges to zero exponentially fast as $n\rightarrow\infty$.
+It is also worth noting that the second part of Theorem 2.6 implies that, under (H1) and if (H2) holds for the discrete metric (that is with $\kappa=\varphi_{0}$), then, for any $V$-semi-distance $\phi$, the $(\varphi_{V},\phi)$-Dobrushin contraction coefficient $\beta_{\varphi_{V},\phi}(P_{n})$ converges to zero exponentially fast as $n\rightarrow\infty$.
 2.6.1 Application 1: Wasserstein exponential contraction estimates
@@
 However, this result further assumes that there exists a $t^{\prime}>0$ such that $\sup_{x\neq y}\ell_{\kappa}(Q_{t^{\prime}})(x,y)<1$ which, as proved in Lemma 1.1, is equivalent to assuming that $\beta_{\kappa}(Q_{t^{\prime}})<1$.
-In the context of $\mathbb{R}^{n}$-valued diffusions, an exponential contraction estimate is obtained in [30], assuming the existence of a Lyapunov function $V$ satisfying the geometric drift condition (2.27) and some additional key growth conditions (see Assumptions 2.3 in [30]).
+In the context of $\mathbb{R}^{m}$-valued diffusions, an exponential contraction estimate is obtained in [30], assuming the existence of a Lyapunov function $V$ satisfying the geometric drift condition (2.27) and some additional key growth conditions (see Assumptions 2.3 in [30]).
 Specifically, Theorem 2.1 in [30] shows that $\beta_{\phi}(Q_{t})\leq\lambda^{t}$ for some constant $\lambda\in]0,1[$, where $\phi(x,y)=\varpi(\|x-y\|)+\varphi_{\rho}(x,y)$ for some judiciously chosen bounded non-decreasing concave continuous functions $\varpi$ and constant $\rho>0$.
@@
 3.1 Some illustrative examples
-3.1.1 Markov chains on the unit open interval with bounded densities
-We let $S=]0,1[$, $\psi_{e}(x,y)=|x-y|$ be the Euclidean metric and, for any $\iota\in]0,1]$ we let
-$\psi_{\iota}(x,y):=\left|\frac{1}{x}-\frac{1}{y}\right|^{\iota}+\left|\frac{1}{1-x}-\frac{1}{1-y}\right|^{\iota}.$ (3.3)
-We note that $(S,\psi_{\iota})$ is a complete and separable metric space for any $\iota\in]0,1]$ while the space $(S,\psi_{e})$ is separable but not complete.
-Next, given a probability measure $\nu(dy)=g(y)dy\in\mathcal{P}(S)$ having a bounded density $g$ and a Markov transition $Q(x,dy)=q(x,y)dy$ on ]0,1[]0,1[ such that $\|q\|<\infty$, we let
-$\begin{array}[]{l}P(x,dy)=x\,Q(x,dy)+(1-x)\nu(dy)\end{array}$ (3.4)
-and, for some $\iota\in]0,1[$, we define the continuous function
+3.1.1 Markov chains on the unit open interval
+We let $S=]0,1[$, $\psi_{e}(x,y)=|x-y|$ be the Euclidean metric, $\iota\in]0,1[$,
+$\psi_{\iota}(x,y):=\left|\frac{1}{x}-\frac{1}{y}\right|^{\iota}+\left|\frac{1}{1-x}-\frac{1}{1-y}\right|^{\iota}$ (3.3)
+and we define the continuous function
 $W_{\iota}(x)=\frac{1}{x^{\iota}}+\frac{1}{(1-x)^{\iota}}.$ (3.4)
-Noting that $W_{\iota}\in{\cal B}_{\infty}(S)$ while $\|P(W_{\iota})\|\leq c:=2(\|q\|+\|g\|)/(1-\iota)<\infty$, it follows that (3.1) holds with $\Theta=1/W_{\iota}$ and thus, by (3.2), condition (H1) holds with $V=W_{\iota}$ (observe that $W_{\iota}$ is bounded below away from zero).
-Assume first that the two probability distributions $\nu(dy)$ and $Q(x,dy)$ have disjoint supports for all $x\in S$.
-In this scenario, we have $\|\delta_{x}P-\delta_{0}P\|_{\tiny tv}=x$ implying that $\beta_{\varphi_{0}}(P)=1$ and therefore (H2) does not hold with the discrete metric, that is with $\kappa=\varphi_{0}$.
-Assume now that, for all $x\in S$, the two probability distributions $\nu(dy)$ and $Q(x,dy)$ have $S$ as support, and assume also that their densities are continuous and strictly positive.
-In this scenario, (H2) holds with $\kappa=\varphi_{0}$ by Lemma 2.1, and Theorem 2.6 therefore ensures that the Kantorovitch semi-distance $D_{\phi}(\mu_{1}P_{n},\mu_{2}P_{n})$ vanishes exponentially quickly as $n\rightarrow\infty$ for any $V$-semi-distance $\phi$.
+We note that $(S,\psi_{\iota})$ is a complete and separable metric space, while the space $(S,\psi_{e})$ is separable but not complete.
+We now let $\nu(dy)=g(y)dy\in\mathcal{P}(S)$ and $\mu(dy)=h(y)dy\in\mathcal{P}(S)$ be probability measures with bounded densities, i.e. such that $\|g\|+\|h\|<\infty$, and we assume first that
+$\begin{array}[]{l}P(x,dy)=x\,\mu(dy)+(1-x)\nu(dy).\end{array}$ (3.5)
+Noting that $W_{\iota}\in{\cal B}_{\infty}(S)$ while $\|P(W_{\iota})\|\leq c:=2(\|g\|+\|h\|)/(1-\iota)<\infty$, it follows that (3.1) holds with $\Theta=1/W_{\iota}$ and thus, by (3.2), condition (H1) holds with $V=W_{\iota}$ (observe that $W_{\iota}$ is bounded below away from zero).
+We further assume that the two probability distributions $\nu$ and $\mu$ have disjoint supports. In this situation, for any $x_{1},x_{2}\in S$ we have
+$\|\delta_{x_{1}}P-\delta_{x_{2}}P\|_{\tiny tv}=|x_{1}-x_{2}|~\|\mu-\nu\|_{\tiny tv}=|x_{1}-x_{2}|$
+implying that $\beta_{\varphi_{0}}(P)=1$. However, noting that,
+$W_{\iota}\leq r\implies P(x,dy)\geq(1-\epsilon(r))\nu(dy),\quad\epsilon(r):=\sup_{W_{\iota}(x)\leq r}x\in]0,1[$
+it follows that the local minorization condition (2.2) holds for any $r_{0}>0$ and with $\nu_{r}=\nu$.
+Hence, condition (H2) is met with the discrete metric, that is with $\kappa=\varphi_{0}$.
+In this context, by Theorem 2.6, for any $V$-semi-distance $\phi$ the Kantorovitch semi-distance $D_{\phi}(\mu_{1}P_{n},\mu_{2}P_{n})$ converges to zero exponentially fast as $n\rightarrow\infty$.
 In particular, noting that the complete metric $\psi_{S}=\psi_{\iota}$ is a $V$-semi-distance function, Theorem 2.6 gives the following exponential contraction estimate for the Wasserstein 1-distance:
 $\mathbb{W}_{\psi_{S}}\left(\mu_{1}P_{n},\mu_{2}P_{n}\right)\leq c\lambda^{n}\|\mu_{1}-\mu_{2}\|_{V},\quad\forall n\geq 0,\quad\forall\mu_{1},\mu_{2}\in\mathcal{P}_{V}(S).$ (3.6)
-Lastly, following an example proposed in [24], Section 2.1, we assume that
+Following an example proposed in [24], Section 2.1, we now assume that
 $P(x,dy):=\frac{1}{2x}~1_{]0,x]}(y)\,dy+\frac{1}{2(1-x)}~1_{]x,1[}(y)\,dy$
 so that the Markov transition $P$ has an unbounded density.
-In this context, we let $\iota\in]0,1/2[$ and show in Section 4.7 that there exist some parameters $\epsilon\in[0,1[$ and $c\geq 0$ such that
+In this situation, we impose that $\iota\in]0,1/2[$ and show in Section 4.7 that there exist some parameters $\epsilon\in[0,1[$ and $c\geq 0$ such that
 $P(W_{\iota})(x)\leq\epsilon\,W_{\iota}(x)+c$ (3.7)
@@
 3.2 Markov chains on domains with a boundary
-In this subsection we consider a Markov transition $P(x,dy)=p(x,y)dy$ with a uniformly bounded, strictly positive and continuous density $p(x,y)$ on a open and connected set $S\subset\mathbb{R}^{n}$ having a non-empty boundary $\partial S$.
+In this subsection we consider a Markov transition $P(x,dy)=p(x,y)dy$ with a uniformly bounded, strictly positive and continuous density $p(x,y)$ on a open and connected set $S\subset\mathbb{R}^{m}$ having a non-empty boundary $\partial S$.
 We define the continuous function
-$d_{n}(x,\partial S):=\inf{\left\{\|x-y\|~:~y\in\partial S\right\}}$
-and, for some $\iota\in]0,1]$, we equip $S$ with the metric
-$\psi_{S}(x,y):=\left|\frac{1}{d_{n}(x,\partial S)}-\frac{1}{d_{n}(y,\partial S)}\right|^{\iota}+\|x-y\|^{\iota}.$ (3.10)
+$d_{m}(x,\partial S):=\inf{\left\{\|x-y\|~:~y\in\partial S\right\}}$
+and, for some $\iota\in]0,1[$, we equip $S$ with the metric
+$\psi_{S}(x,y):=\left|\frac{1}{d_{m}(x,\partial S)}-\frac{1}{d_{m}(y,\partial S)}\right|^{\iota}+\|x-y\|^{\iota}.$ (3.10)
 It can be shown that $(S,\psi_{S})$ is a complete and separable metric space.
 3.2.1 Bounded domains with a locally Lipschitz boundary
-Assume that $S$ is bounded and has a locally Lipschitz boundary $\partial S$, and let $W(x)=d_{n}(x,\partial S)^{-\iota}+\|x\|$.
+Assume that $S$ is bounded and has a locally Lipschitz boundary $\partial S$, and let $W(x)=d_{m}(x,\partial S)^{-\iota}+\|x\|^{\iota}$.
 In this scenario, we clearly have $W\in{\cal B}_{\infty}(S)$ and, since the function $W$ is integrable over the set $S$ (c.f.
@@
 On the other hand, since $P$ is assumed to have a strictly positive and continuous density, condition (H2) holds with $\kappa=\varphi_{0}$ by Lemma 2.1.
-By Theorem 2.6, it follows that for any $V$-semi-distance $\phi$ the Kantorovitch semi-distance $D_{\phi}(\mu_{1}P_{n},\mu_{2}P_{n})$ converges to zero exponentially quickly as $n\rightarrow\infty$.
+By Theorem 2.6, it follows that for any $V$-semi-distance $\phi$ the Kantorovitch semi-distance $D_{\phi}(\mu_{1}P_{n},\mu_{2}P_{n})$ converges to zero exponentially fast as $n\rightarrow\infty$.
 In particular, one can check that the metric $\psi_{S}$ defined in (3.10) is the $V$-semi-distance function, and thus Theorem 2.6 gives the following exponential contraction estimate for the Wasserstein 1-distance:
@@
 Assume that $S$ is not necessarily bounded but has a locally Lipschitz bounded boundary $\partial S$.
-Assume also that there exists a function $W_{1}\in{\cal B}_{\infty}(\mathbb{R}^{n})$ such that $\sup_{x\in S}\|x\|^{\iota}/W_{1}(x)<\infty$ and such that $P(W_{1})\leq\epsilon_{1}W_{1}+c_{1}$ for some parameters $\epsilon_{1}\in[0,1[$ and $c_{1}>0$.
+Assume also that there exists a function $W_{1}\in{\cal B}_{\infty}(\mathbb{R}^{m})$ such that $\sup_{x\in S}\|x\|^{\iota}/W_{1}(x)<\infty$ and such that $P(W_{1})\leq\epsilon_{1}W_{1}+c_{1}$ for some parameters $\epsilon_{1}\in[0,1[$ and $c_{1}>0$.
 In this scenario, condition (H1) holds with $V=1+W_{1}$ but, for this definition of $V$, the metric $\psi_{S}$ defined in (3.10) may not be a $V$-semi-distance.
-To solve this problem we let $W_{2}(x)=d_{n}(x,\partial S)^{-\iota}\in{\cal B}_{\infty}(S)$ and $c_{2}=\|P(W_{2})\|\leq 1+\|P(W_{2}1_{W_{2}\geq 1})\|$.
-Using the fact that $\{x\in S:W_{2}(x)\geq 1\}=\{x\in S:d_{n}(x,\partial S)\leq 1\}$ with $\partial S$ a bounded set, and recalling that the density $p(x,y)$ is assumed to be uniformly bounded, by using a similar argument as in the proof of Lemma 5.2 in [4] we can check that $\|P(W_{2}1_{W_{2}\geq 1})\|<\infty$.
+To solve this problem we let $W_{2}(x)=d_{m}(x,\partial S)^{-\iota}$ and $c_{2}=\|P(W_{2})\|\leq 1+\|P(W_{2}1_{W_{2}\geq 1})\|$.
+Using the fact that $\{x\in S:W_{2}(x)\geq 1\}=\{x\in S:d_{m}(x,\partial S)\leq 1\}$ with $\partial S$ a bounded set, and recalling that the density $p(x,y)$ is assumed to be uniformly bounded, by using a similar argument as in the proof of Lemma 5.2 in [4] we can check that $\|P(W_{2}1_{W_{2}\geq 1})\|<\infty$.
 This implies that $c_{2}<\infty$ and thus, since
@@
 Example 3.1.
-Assume that $(g,h)$ and $(m_{g},m_{h})$ are continuous functions. In addition, assume that there exist some parameters $a_{1},b_{1}\in\mathbb{R}$, $a_{2},b_{2}>0$ and $q\geq p>0$ such that
+Assume that $(g,h)$ and $(m_{g},m_{h})$ are continuous functions on $S=\mathbb{R}^{d}$.
+In addition, assume that there exist some parameters $a_{1},b_{1}\in\mathbb{R}$, $a_{2},b_{2}>0$ and $q\geq p>0$ such that
 $m(x,y)\leq a_{1}+a_{2}~(\|x\|^{p}+\|y\|^{p})\quad\mbox{and}\quad g(x)\wedge h(x)\geq b_{1}+b_{2}~\|x\|^{q}.$
@@
 We conclude that $P(W)/W\leq\Theta\in{\cal B}_{0}(S)$ with the functions
-$W(x):=\frac{1}{2}~\exp{(\delta\|x\|)}\quad\mbox{\rm and}\quad\Theta(x):=a_{1}~\exp{\left(-\delta(1-\lambda_{0})~\|x\|\right)}$
+$W(x):=\exp{(\delta\|x\|)}\quad\mbox{\rm and}\quad\Theta(x):=a_{1}~\exp{\left(-\delta(1-\lambda_{0})~\|x\|\right)}$
 and thus, by (3.2), condition (H1) holds with $V=W\in{\cal B}_{\infty}(S)$.
-Assume now that $\mathbb{E}\left(\|G_{Z}(x)\|^{p}\right)<\infty$ for some $p\geq 1$.
+Assume now that $\sup_{x\in S}\mathbb{E}\left(\|G_{Z}(x)\|^{p}\right)<\infty$ for some $p\geq 1$.
 In this scenario, letting $c_{\lambda,p}=c_{\lambda}+\sup_{x\in S}\mathbb{E}\left(\|G_{Z}(x)\|^{p}\right)^{1/p}$ we observe that
@@
 3.4.3 Example 1: Random iterated functions with local dissipative conditions
-Consider the functions $(F,F_{z},G_{z})$ defined in Section 3.4.2 with ${\cal Z}=\mathbb{R}^{n}=S$ and
+Consider the functions $(F,F_{z},G_{z})$ defined in Section 3.4.2 with ${\cal Z}=\mathbb{R}^{m}=S$ and
 $F(x)=x+b(x)~h\quad\mbox{\rm and}\quad G_{z}(x)=z$
@@
 This shows that condition (3.26) is met with $x_{0}=0$ and the results of the previous subsection can be used to obtained a function $V\in{\cal B}_{\infty}(S)$ with compact level sets such that (H1) holds.
-Specifically, if $Z$ is a Gaussian random variable it follows from the calculations in Section 3.4.2 that (H1) is met when $V=(1/2)\exp(\delta\|x\|)$ for some $\delta>0$.
+Specifically, if $Z$ is a Gaussian random variable it follows from the calculations in Section 3.4.2 that (H1) is met when $V=\exp(\delta\|x\|)$ for some $\delta>0$.
 If we further assume that the function $b$ is continuous, it follows from the discussion of Section 3.4.1 that condition (H2) is met with $\kappa=\varphi_{0}$.
@@
 Let $F:S\rightarrow S$ be a mapping such that (3.26) holds for some $x_{0}\in S$ and parameters $\lambda(x_{0})\in]0,1[$ and $c(x_{0})\geq 0$.
-Next, we let $\delta>0$, $W(x)=\frac{1}{2}\exp{(\delta\|x\|)}$ and assume that there is an $r_{0}>0$ such that, for all $r\geq r_{0}$, there exists an $\alpha(r)\in]0,1[$ such that
+Next, we let $\delta>0$, $W(x)=\exp{(\delta\|x\|)}$ and assume that there is an $r_{0}>0$ such that, for all $r\geq r_{0}$, there exists an $\alpha(r)\in]0,1[$ such that
 $W(x)+W(y)\leq r\implies\|F(x)-F(y)\|\leq(1-\alpha(r))\|x-y\|.$ (3.29)
@@
 3.5.1 A general diffusion model
-We let $S=\mathbb{R}^{n}$, $\psi_{S}(x,y)=\|x-y\|$ and we consider below the continuous time Markov semigroup $(Q_{t})_{t\in]0,\infty[}$ of a diffusion flow $y\in S\mapsto X_{t}(y)$ defined by
+We let $S=\mathbb{R}^{m}$, $\psi_{S}(x,y)=\|x-y\|$ and we consider below the continuous time Markov semigroup $(Q_{t})_{t\in]0,\infty[}$ of a diffusion flow $y\in S\mapsto X_{t}(y)$ defined by
 $dX_{t}(y)=b(X_{t}(y))\,dt+\tau(X_{t}(y))\,dB_{t}\quad\mbox{\rm with}\quad X_{0}(y)=y.$ (3.30)
-In the above display, $B_{t}$ is a $n_{1}$-dimensional Brownian motion starting at the origin for some $n\geq 1$ and $b$ is a differentiable drift function from $\mathbb{R}^{n}$ into itself satisfying, for some constants $a>0$ and $c\geq 0$,
+In the above display, $B_{t}$ is an $m_{1}$-dimensional Brownian motion starting at the origin and $b$ is a differentiable drift function from $\mathbb{R}^{m}$ into itself satisfying, for some constants $a>0$ and $c\geq 0$,
 $x^{\top}b(x)\leq-a\,\|x\|^{2}+c,\quad\forall x\in S.$ (3.31)
-Moreover, $\tau$ stands for some $(n\times n_{1})$-matrix valued continuous function such that, for all $x\in S$, the matrix $\tau(x)\tau(x)^{\top}$ is invertible and the map $x\mapsto\|\tau(x)\|/\|x\|$ vanishes at infinity.
+Moreover, $\tau$ stands for some $(m\times m_{1})$-matrix valued continuous function such that, for all $x\in S$, the matrix $\tau(x)\tau(x)^{\top}$ is invertible and the map $x\mapsto\|\tau(x)\|/\|x\|$ vanishes at infinity.
 In what follows, for any $x\in S$ we let
-$\|\tau(x)\|_{F}^{2}:=\mbox{\rm Tr}(\tau(x)\tau(x)^{\top})\leq n_{1}\|\tau(x)\|_{2}^{2}\quad\text{ with }\|\tau(x)\|_{2}^{2}=\lambda_{\max}(\tau(x)\tau(x)^{\top})$
+$\|\tau(x)\|_{F}^{2}:=\mbox{\rm Tr}(\tau(x)\tau(x)^{\top})\leq m_{1}\|\tau(x)\|_{2}^{2}\quad\text{ with }\|\tau(x)\|_{2}^{2}=\lambda_{\max}(\tau(x)\tau(x)^{\top})$
 and for all $p\geq 1$ we let $W^{p}(x):=\|x\|^{p}$.
@@
 3.5.2 The overdamped Langevin diffusion model
-An important class of diffusion processes arising in statistical physics and applied probability is the overdamped Langevin diffusion flow $X_{t}(y)$, defined as in (3.30) with $(n_{1},\tau(x))=(n,\sigma I)$ for some $\sigma>0$ and the drift function $b(x)=-\gamma\nabla U(x)$ associated with some parameter $\gamma>0$ and smooth potential function $U$ satisfying the following condition
+An important class of diffusion processes arising in statistical physics and applied probability is the overdamped Langevin diffusion flow $X_{t}(y)$, defined as in (3.30) with $(m_{1},\tau(x))=(m,\sigma I)$ for some $\sigma>0$ and the drift function $b(x)=-\gamma\nabla U(x)$ associated with some parameter $\gamma>0$ and smooth potential function $U$ satisfying the following condition
 $\nabla^{2}U(x)\geq\lambda~1_{\|x\|\geq r}~I-c~1_{\|x\|<r}~I\quad\mbox{for some $\lambda>0$, $c\geq 0$ and $r\geq 0$.}$ (3.33)
@@
 Following the proof of Proposition 2.16 in [4], for any $\delta>0$ and $h>0$ there exist some constants $\delta_{h}>0$ and $c_{h}>0$ such that
-$W(x):=(1/2)\exp(\delta\|x\|)\Longrightarrow\sup_{0\leq h^{\prime}\leq 1}\|\widetilde{Q}_{h^{\prime}}(W)/W\|<\infty\text{ and }\widetilde{Q}_{h}(W)/W\leq c_{h}/W^{\delta_{h}}.$ (3.34)
+$W(x):=\exp(\delta\|x\|)\Longrightarrow\sup_{0\leq h^{\prime}\leq 1}\|\widetilde{Q}_{h^{\prime}}(W)/W\|<\infty\text{ and }\widetilde{Q}_{h}(W)/W\leq c_{h}/W^{\delta_{h}}.$ (3.34)
 On the other hand, using the stochastic interpolation formula (cf. Theorem 1.2 in [23]) we check the almost sure estimate uniform estimate
@@
 $D_{\phi_{k}}(\mu_{1}P,\mu_{2}P)$ $\leq(\pi Q_{k})(\phi_{k})\leq\ell_{\psi,\phi}^{\star}(P)\pi(\psi)=\ell_{\psi,\phi}^{\star}(P)D_{\psi}(\mu_{1},\mu_{2})$ (4.8)
-showing that it $D_{\psi}(\mu_{1},\mu_{2})>0$ then
+showing that if $D_{\psi}(\mu_{1},\mu_{2})>0$ then
 $\frac{D_{\phi_{k}}(\mu_{1}P,\mu_{2}P)}{D_{\psi}(\mu_{1},\mu_{2})}\leq\ell_{\psi,\phi}^{\star}(P).$
@@
 4.4.1 Proof of (2.14)
-Let $x,y\in S$ be such that $x\neq y$ and $r>r_{\epsilon}\vee r_{0}$ be fixed in what follows. Then, under the assumption that (H2) holds with the semi-distance $\kappa=\varphi_{0}$, we have
-$\|\delta_{x}P-\delta_{y}P\|_{\tiny tv}\leq 1-\alpha(r)<1.$ (4.12)
-In what follows the functions $(\varpi_{\rho},\varphi_{\rho})$ indexed by $\rho>0$ are as defined in (2.3) and (2.4).
+Let $x,y\in S$ be such that $x\neq y$ and $r>r_{\epsilon}\vee r_{0}$ be fixed in what follows.
+In addition, below the functions $(\varpi_{\rho},\varphi_{\rho})$ indexed by $\rho>0$ are as defined in (2.3) and (2.4).
 Assume first that $\varpi_{V}(x,y)\geq r$. In this scenario, for all $\rho>0$ we have
-$\begin{split}\ell_{\varphi_{\rho}}(P)(x,y)&\leq\ell_{\varpi_{\rho}}(P)(x,y)=\frac{1+\rho P(V)(x)+\rho P(V)(y)}{1+\rho\varpi_{V}(x,y)}\\ &\leq\frac{1+\rho\big(1+\epsilon\varpi_{V}(x,y)\big)}{1+\rho\varpi_{V}(x,y)}\\ &=\epsilon+\frac{1+\rho-\epsilon}{1+\rho r}=1-(1-\epsilon)~\frac{\rho r}{1+\rho r}~\left(1-\frac{r_{\epsilon}}{r}\right)\end{split}$ (4.13)
-where the first inequality uses the assumption that (H1) holds with $c=1/2$ and the last inequality uses (4.12). Simple calculations show that
+$\begin{split}\ell_{\varphi_{\rho}}(P)(x,y)&\leq\ell_{\varpi_{\rho}}(P)(x,y)=\frac{1+\rho P(V)(x)+\rho P(V)(y)}{1+\rho\varpi_{V}(x,y)}\\ &\leq\frac{1+\rho\big(1+\epsilon\varpi_{V}(x,y)\big)}{1+\rho\varpi_{V}(x,y)}\\ &\leq\epsilon+\frac{1+\rho-\epsilon}{1+\rho r}=1-(1-\epsilon)~\frac{\rho r}{1+\rho r}~\left(1-\frac{r_{\epsilon}}{r}\right)\end{split}$ (4.12)
+where the second inequality uses the assumption that (H1) holds with $c=1/2$. Simple calculations show that
 $(1-\epsilon)\frac{r\rho_{\epsilon}(r)}{1+r\rho_{\epsilon}(r)}\left(1-\frac{r_{\epsilon}}{r}\right)$ $=~\frac{\alpha(r)}{2}~\frac{1-\epsilon}{(1+\epsilon)+\alpha(r)/2}~\left(1-\frac{r_{\epsilon}}{r}\right)\geq~\frac{\alpha(r)}{2}~\delta_{\epsilon}(r)$